### v1

Abstract
Can the causal effects estimated in experiment be generalized to real-world scenarios? This question lies at the heart of social science studies. External validity primarily assesses whether experimental effects persist across different settings, implicitly presuming the experiment’s ecological validity—that is, the consistency of experimental effects with their real-life counterparts. However, we argue that this presumed consistency may not always hold, especially in experiments involving multi-dimensional decision processes, such as conjoint experiments. We introduce a formal model to elucidate how attention and salience effects lead to three types of inconsistencies between experimental findings and real-world phenomena: amplified effect magnitude, effect sign reversal, and effect importance reversal. We derive testable hypotheses from each theoretical outcome and test these hypotheses using data from various existing conjoint experiments and our own experiments. Drawing on our theoretical framework, we propose several recommendations for experimental design aimed at enhancing the generalizability of survey experiment findings.
Keywords: Conjoint Experiment, Survey Experiment, Ecological Validity, External Validity, Generalizability, Attention, Salience.
[…]
Introduction
The randomized experiment is often hailed as the gold standard for identifying and estimating causal effects within the social sciences. One popular such method is conjoint experiment, which is frequently employed by researchers to examine how individual preferences for candidates or immigrants are influenced by characteristics such as gender, race, experience, and skills (Eggers, Vivyan and Wagner, 2018; Carnes and Lupu, 2016; Sanbonmatsu, 2002). However, the interest of researchers extends well beyond the mere observation of causal effects within an experimental setting. Especially, we ask whether these experimental estimates accurately mirror real-world effects. For instance, in a conjoint survey experiment that reveals a preference for female candidates with an average causal effect of approximately 2%, researchers hope that this finding corresponds to an actual increase in preference for female candidates in real-world scenarios. When such a correspondence is established, we regard the experimental effect as being consistent with the real-world effect. This consistency, which establishes the ecological validity of experiment, serves as a crucial concern of social science studies.
Ecological validity represents a critical aspect of external validity, distinct from internal validity, which pertains to the validity of conclusions within the experimental setting (Kihlstrom 2021). As depicted in Figure 1, external validity primarily concerns the generalization of results from the experimental domain to real-world contexts with varied populations, treatments, outcomes, and contexts (Cook, Campbell and Shadish 2002). A fundamental premise is that the effects observed in an experiment should be consistent with real-life effects. If the effects estimated in the experiment do not align with those experienced by the same individuals in the real world, it undermines the rationale for discussing the generalizability of these results to other populations, treatments, outcomes, or contexts.
Figure 1: Experimental Validity: Ecological validity as a premise or a special subset of External Validity
We demonstrate that experiments involving multi-dimensional decision-making processes, such as the widely used vignette and conjoint survey experiments with multiple attributes manipulated at random, frequently fail to achieve ecological validity. This failure raises substantial concerns about our ability to generalize experimental insights to other real-life scenarios effectively. Within the scope of this paper, we specifically examine three types of consistency: effect magnitude, effect direction, and attribute importance. The Consistency of Effect Magnitude asserts that the observed size of an experimental effect should mirror the magnitude (and the sign) observed in real-world phenomena. This criterion, while offering the deepest insights into real-world applicability, stands as the most challenging to fulfill. The Consistency of Effect Direction posits that the sign (positive or negative) of an experimental effect should match its real-world counterpart. Achieving magnitude consistency ensures direction consistency by default, rendering the latter a less rigorous yet still crucial requirement for real-world insights. Lastly, Consistency of Attribute Importance emphasizes the need for experimental outcomes to truthfully represent the relative magnitude of different attributes. This aspect is particularly valuable for researchers aiming to identify which attributes most significantly influence variations in outcomes, marking it as another derivative of magnitude consistency.
It is not imperative for an experiment to meet all three consistency criteria simultaneously; rather, the requisite consistency should align with the experiment’s objectives. For instance, program evaluations conducted by policymakers, who require information on effect magnitude to assess overall welfare changes, necessitate experiments to meet the criterion of magnitude consistency. Conversely, when experiments are designed to test predictions derived from formal theory, often only direction consistency is essential. And, in some cases, the intentionally amplified causal effect can even help researchers to reach the statistical power. Similarly, attribute importance consistency is sufficient when the goal is to understand attribute influences and to predict outcomes [^3] In machine learning exercises, for example, predictors that account for the majority of variation are assigned greater importance and have a more significant impact on predictions..
We develop a formal theory to elucidate why experiments involving multi-dimensional decision-making frequently fall short of meeting the three consistency criteria. At the heart of our theoretical framework lies the interplay between attention and salience. Conventional decision-making theories have often posited that individuals possess unlimited attention and exhibit consistent salience across contexts (i.e. they take account for all possible factors and the weight for each factor is independent or unchanged in different situations). While these assumptions yield straightforward theoretical predictions, they have been increasingly called into question due to numerous observed anomalies and paradoxes. Notably, the Allais (1953) and Ellsberg (1961) paradoxes serve as prominent examples that challenge these traditional views. The former one shows that individual frequently violate the independence axiom of expected utility theory and the latter one shows that people averse ambiguity and their decisions are inconsistent with subjective expected utility theory. While the bulk of existing research has focused on refining decision theory to account for these discrepancies, our work pioneers the application of a behavioral theoretical approach to systematically address these inconsistencies within experimental research methodologies.
Our primary formal finding underscores that experimental effects involving multi-dimensional decision-making processes are notably exaggerated due to the phenomena of limited attention, which is also observed by Barabas and Jerit (2010) in the context of survey experiments. Attention effects posit that the consideration set within an experimental context diverges significantly from that in real-world scenarios. The limitation and transience of human attention have long been recognized by social scientists. For example, consumer choices may favor option $y$ in the presence of $x$, but this preference can reverse when $x$ is not immediately apparent to the consumer (Masatlioglu, Nakajima and Ozbay, 2016). Chetty’s field experiments demonstrate that consumers tend to under-react to taxes when tax-inclusive pricing is not prominently displayed, leading to an 8 percent decrease in demand when it is (Chetty, Looney and Kroft, 2009). We extend these empirical findings by providing a theoretical framework that outlines sufficient conditions under which this phenomenon occurs. In doing so, we formalize the idea that the experimental setting can inadvertently direct and intensify individual attention to certain attributes, resulting in biased estimates relative to the real-world behavior.
To elucidate the inconsistencies in the direction and relative importance of experimental effects, we delve into the salience effect. The main intuition is that in a complex decision-making environment, the salience effect can induce correlation between otherwise independent attributes, significantly altering the weight placed on certain attributes and potentially reversing decision outcomes. To precisely model this salience effect, we draw on the psychologically grounded model of choice proposed by Bordalo, Gennaioli and Shleifer (2012, 2013), which posits that salience is primarily driven by the perceived intensity of each attribute.
We derive several corresponding testable hypotheses from our theoretical results. For example, a key hypothesis is that the observed effect diminishes or even reverses as the number of attributes increases. To empirically test these hypotheses, we analyze data from multiple existing studies and conduct additional, pre-registered conjoint experiments  [^4] The study is pre-registered and approved by authors’ Institutional Review Board; More details are in the SI D.. The experimental findings consistently support our hypotheses, thereby reinforcing our theoretical framework. We do not claim that our proposed theory is the only reason that generates those inconsistency. However, from our theory, we can derive some specific hypotheses that may not straightforwardly derived from other theories. We believe that the positive results from testing those hypotheses could further support our theory. In light of mechanisms driving inconsistencies identified in our theory, we propose several experimental design recommendations in the discussion section aimed at alleviating these issues.
Our research intersects with and contributes to several vital streams of literature. At the core of our study is a pivotal inquiry into the conjoint experiment literature, a timely and critical exploration considering the burgeoning body of research on conjoint analysis since the seminal work of Hainmueller, Hopkins and Yamamoto (2014). Subsequent investigations have primarily honed in on statistical concerns—such as external validity (De la Cuesta, Egami and Imai, 2022), measurement errors (Clayton et al., 2023), power analysis (Stefanelli and Lukac, 2020), survey satisficing (Bansak et al., 2018, 2021), social desirability bias (Horiuchi, Markovich and Yamamoto, 2022), hypotheses testing (Ham, Imai and Janson, 2022), and multiple testing (Horiuchi, Markovich and Yamamoto, 2022)—yet, the fundamental issue of real-world consistency remains underexplored. Hainmueller, Hangartner and Yamamoto (2014) empirically find that survey experiments have high ecological validity [^5] In their paper, they primarily describe it as external validity. Actually, they demonstrate both external validity (comparing different individuals) and ecological validity (comparing experimental effects and real-world effects).. One main reason is that their experiment successfully eliminates the attention effect so that the consideration set of the experiment is almost the same to the real-world one. Our findings provide a theoretical explanation to understand conditions under which conjoint experiments are reliable.
Second, our work engages with the emerging literature on Theoretical Implication of Empirical Methods (TIEM) (Slough and Tyson, 2023; De Mesquita and Tyson, 2020; Slough, 2023; Fu and Slough, 2023), which critically assesses empirical methodologies through the lens of formal theory. This approach, which contrasts with the Empirical Implications of Theoretical Models (EITM), enables a deeper understanding of the limitations and potential inconsistencies within empirical research methods. In the context of conjoint experiments, our application of decision theory and behavioral economics illuminates foundational discrepancies between experimental effects and their real-world counterparts, extending beyond the aggregate reference inconsistencies identified by Abramson, Koçak and Magazinnik (2019).
Third, we contribute to the discussion on external validity, which is a core concern in experimental research, especially in lab and survey experiments (Barabas and Jerit 2010; List and Levitt 2005; Gaines, Kuklinski and Quirk 2007). Slough and Tyson (2023, 2024) establish a formal framework to clarify the external validity. They identify conditions when different studies are target-equivalence and target-congruence. They particularly emphasize whether studies can be harmonized along treatment, control, and the measurement of the outcome. Egami and Hartman (2023) also develop a unified framework to discuss four types of external validity, including population, contexts, treatments, and outcomes, as proposed by Cook, Campbell and Shadish (2002). Several estimators are then proposed to conduct effect generalization and sign generalization. Most studies primarily focus on whether the results can be generalized to a new population (Mullinix et al. 2015; Huang 2022). Our paper focuses on ecological validity, which compares the experimental effect and real-world effect even for the same individuals, treatments, and outcomes. Many factors can affect ecological validity, such as the famous Hawthorne effect (Adair, 1984). We provide a new explanation to analyze ecological validity for experiments involving multi-dimensional decision processes.
Lastly, we broaden the application of salience theory beyond its traditional confines in political science and economics, where its implications for election, party competition, agenda-setting, and judicial biases have been well-documented (Dragu and Fan, 2016; Riker, Riker and Riker, 1986; Ascencio and Gibilisco, 2015; Guthrie, Rachlinski and Wistrich, 2000; Viscusi, 2001). By integrating salience theory into experimental design and applying the framework developed by Bordalo, Gennaioli and Shleifer (2012, 2013, 2016), we highlight salience as a pivotal concern in research methodologies, extending its relevance and applicability.
1 Theoretical Model of Experiment
A decision maker (DM) evaluates alternatives/profiles within a set $A=\{A_{1},A_{2},...,A_{j}\}$. Each alternative is characterized by $K$ attributes, denoted by the vector $X=(X_{1},...,X_{K})$. In a typical candidate-choice conjoint experiment, the DM faces two hypothetical candidates, $A_{1}$ and $A_{2}$. For each candidate, researchers provide several attributes, such as age, gender, and job experience. For DM $i$, the evaluation of each attribute $k$ is denoted by $u_{k}(X_{k},\theta_{i})$, where $\theta_{i}$ represents the DM’s characteristics that may influence individual preferences, including factors like gender, income, education, etc. In the context of multi-dimensional decision-making, the preference for a certain alternative depends on the overall evaluation of its attributes. In other words, the DM assigns a weight $\alpha_{k}$ to each attribute $X_{k}$. We refer to this weight $\alpha$ as “salience.” As we will demonstrate, salience can be easily manipulated or altered by the context. The discussion on how salience changes will be deferred to the section 3. The complete evaluation of alternative $A$ is thus a weighted sum as follows:
$V_{i}(A):=V_{i}(X)=\sum_{k=1}^{K}\alpha_{k}u_{k}(X_{k},\theta_{i})$ (1)
where $\sum_{\alpha_{k}\in\alpha}\alpha_{k}=1$.In the model, we implicitly assume that the evaluation of attribute $X_{k}$ is independent of attribute $X_{k^{\prime}}$ for $k\neq k^{\prime}$. More generally, however, we acknowledge the possibility that they might be correlated. As we will see later, the salience effect can induce correlation among independent attributes. Therefore, the current utility function does not lose generality; rather, it enables us to isolate and understand the mechanism of the salience effect.
Within the context of the above model, we define the Individual Treatment Effect (ITE) for attribute $X_{1}$. Consistent with standard practice, the causal effects are understood in terms of potential outcomes in a hypothetical scenario. We use Table 1 to illustrate the concept. Suppose we aim to define the ITE for attribute $X_{1}$ when it changes from $x_{1}$ to $\tilde{x}_{1}$, while holding the other attributes constant. This involves considering two hypothetical scenarios.
$World1$
$A_{1}$ $A_{2}$
$X_{1}$ $x_{1}$ $x^{\prime}_{1}$
$X_{2}$ $x_{2}$ $x^{\prime}_{2}$
$X_{3}$ $x_{3}$ $x^{\prime}_{3}$
$World2$
$\tilde{A}_{1}$ $A_{2}$
$\tilde{x}_{1}$ $x^{\prime}_{1}$
$x_{2}$ $x^{\prime}_{2}$
$x_{3}$ $x^{\prime}_{3}$
Table 1: Illustration of Multi-dimensional Decisions. $A_{j}$ is the alternatives and $X_{k}$ is the attribute.
In the first scenario (World 1), DMs are presented with a choice between two candidates, $A_{1}$ and $A_{2}$. Each candidate is characterized by three attributes: $X_{1},X_{2}$, and $X_{3}$. The realized values of these attributes for $A_{1}$ and $A_{2}$ are the vectors $(x_{1},x_{2},x_{3})$ and $(x^{\prime}_{1},x^{\prime}_{2},x^{\prime}_{3})$, respectively. For example, if $X_{1}$ represents gender, with $x_{1}$ indicating female and $x^{\prime}_{1}$ indicating male, the evaluations for the two candidates in World 1 are $V_{i}(A_{1})$ and $V_{i}(A_{2})$.
In the second scenario (World 2), the DM is presented with the same $A_{2}$, but for $A_{1}$, the attribute $X_{1}$ is realized as $\tilde{x}_{1}$ instead. In other words, all attribute realizations remain the same except for $X_{1}$. We thus use $\tilde{A}_{1}$ to differentiate this modified version of $A_{1}$ from its original. The potential evaluations in World 2 become: $V_{i}(\tilde{A}_{1})$ and $V_{i}(A_{2})$.
In practice, most conjoint experiments measure the forced choice outcome $Y$; that is, whether a candidate is chosen or not, denoted by $1$ or 00. Theoretically, the DM chooses candidate $A_{1}$ over $A_{2}$ because the DM has a higher evaluation for the former. Therefore, $Y$ is a function of the utility difference. The corresponding utility differences in World 1 are $V^{1}_{i}=V_{i}(A_{1})-V_{i}(A_{2})$, and in World 2, they are $V^{2}_{i}=V_{i}(\tilde{A}_{1})-V_{i}(A_{2})$  [^6] Note that $V_{i}(A_{2})$ in the two hypothetical worlds may differ due to the salience effect, which we will discuss in Section 3.. It is evident that the binary choice potential outcome can be defined as $Y_{i}^{j}=\mathbbm{1}[V^{j}_{i}\geq 0]$, where $j=1,2$.
Because the utility difference drives the potential outcome we observe, without loss of generality, throughout the main text, we focus on $V$ rather than $Y$  [^7] For outcomes that are preference scores, our results also hold because the score is also a function of the evaluation. In some survey experiments, the DM may only observe one alternative. In such cases, we can simply set $V_{i}(A_{2})=0$.. Consequently, the individual causal effect for attribute $X_{1}$ when it changes from $x_{1}$ to $\tilde{x}_{1}$, given the other attributes remain constant, is defined as the difference-in-differences $V^{1}_{i}-V^{2}_{i}$.
Definition 1 (ITE).
Given an alternative set $A=\{A_{1},\tilde{A_{1}},A_{2}\}$ characterized by attributes $X=(X_{1},X_{2},...,X_{K})$. Let $X^{A_{1}}=(x_{j},x_{-j})$ represent the attributes of $A_{1}$, where $x_{j}$ is the value of attribute $X_{j}$ for $A_{1}$ and $x_{-j}$ denotes the values of all other attributes except $X_{j}$. Similarly, let $\tilde{X}^{\tilde{A}_{1}}=(\tilde{x}_{j},x_{-j})$ and $X^{A_{2}}=(x^{\prime}_{j},x^{\prime}_{-j})$ represent the attributes of $\tilde{A}_{1}$ and $A_{2}$. For a DM $i$, the individual treatment effect of attribute $X_{j}$ with respect to changing its value from $x_{j}$ to $\tilde{x}_{j}$ is defined as: $ITE^{i}=[V_{i}(X^{A_{1}})-V_{i}(X^{A_{2}})]-[V_{i}(\tilde{X}^{A_{1}})-V_{i}(X^{A_{2}})]$.
This concept is also referred to as the individual component effect (ICE) by Abramson et al. (2023). The formulation aligns with methodologies employed in the conjoint experiment literature, as discussed by Hainmueller, Hopkins and Yamamoto (2014). Given that alternatives are characterized by attributes $X$, we also use the notation $ITE^{i}=[V_{i}(A_{1})-V_{i}(A_{2})]-[V_{i}(\tilde{A}_{1})-V_{i}(A_{2})]$ to denote ITE in this paper. This example of ITE underscores the notion that, within a multi-dimensional decision-making environment, the ITE is contingent not only on the treatment attribute $X_{1}$ but also on other attributes.
It is worth noting that the formal derivation of the ITE hinges on other assumptions: Stability and No Carryover Effects, and No Profile-Order Effects, which are standard in the conjoint experiment literature. A more formal discussion is available in the Supplementary Information (SI) 4. The Marginal Component Effect (MCE) is defined as a weighted sum of ITEs, and the Average Marginal Component Effect (AMCE) represents the expected value of MCE across the distribution of DMs.
Our analytical framework is not confined to conjoint experiments alone; it possesses the flexibility to be applied across a spectrum of (survey) experimental designs. For example, most experiments frequently introduce background information on a given topic, which can be methodically represented through attributes and levels. Until now, our discussion has primarily operated under the presumption that DMs across both treatment and control groups are exposed to an identical array of attributes. Nevertheless, certain designs of survey experiments diverge from this norm, especially in cases where DMs in the control group are not informed about specific attributes that are subjected to manipulation in the treatment group (for a detailed discussion, see Gaines, Kuklinski and Quirk (2007)). One common scenario involves assessing the causal impact of specific information (e.g., about corruption, denoted as $X_{1}$) on individual preferences. In such studies, DMs in both the treatment and control groups might be presented with the same foundational background information, with the distinction that only participants in the treatment group receive details concerning $X_{1}$. This specific experimental setup and its implications for attention effects are further elaborated in the SI B.
2 Attention Effect
Individuals have limited attention spans. Economists have observed that when people make decisions, such as purchasing a computer, they do not consider all possible options (Hausman, 2008). Within the literature, the subset of alternatives that a DM considers when making a decision is referred to as the “consideration set” (Hausman, 2008; Wright and Barbour, 1977). It has been well-argued and demonstrated that due to cognitive limitations, humans cannot allocate sufficient attention to every potential attribute (Stigler, 1961; Jones and Baumgartner, 2005; Chetty, Looney and Kroft, 2009). This indicates that the consideration set is typically smaller than the full set $X$ that may influence a decision. We denote the consideration set in real-world decision-making as $X^{r}\subset X$ and the consideration set in an experimental context as $X^{e}\subset X$.
In experiments, the consideration set of DMs is significantly influenced by the specific context of the experiment. For instance, in a candidate choice experiment, if researchers provide information only on age and gender, respondents’ attention is likely focused exclusively on these two attributes. Thus, the following assumption of limited attention is upheld.
Assumption 1 (Limited Attention).
$X^{e}\subset X^{r}.$
The consideration set utilized in experiments, denoted as $X^{e}$, is often a subset of the real-world consideration set, $X^{r}$, for several reasons. Firstly, researchers seldom possess complete knowledge of which attributes constitute the real-world consideration set $X^{r}$. Secondly, even with such knowledge, it is impractical to present all relevant information about these attributes within the confines of a single survey experiment  [^8] A more general assumption could be that $X^{e}\neq X^{r}$. However, the empirical results in the paper provide little evidence on this assumption..
Without loss of generality, we assume that the real-world consideration set is $X^{r}=\{X_{1},X_{2},...,X_{m}\}$ and the experimental consideration set is $X^{e}=\{X_{1},X_{2},...,X_{k}\}$, such that the cardinality of $X^{r}$, denoted by $m$, is greater than that of $X^{e}$, denoted by $k$ ($|X^{r}|=m>|X^{e}|=k$). In other words, the consideration set for the experiments contains only the first $k$ attributes from the real-world consideration set.
2.1 Inconsistency of Effect Magnitude
Now, let us provide an intuition for how limited attention influences the magnitude of causal effects. Suppose that in the experiment, DMs are asked to make a comparison between two alternatives, $A_{1}$ and $A_{2}$. Let’s say $A_{1}$ has attributes realized as $X^{A_{1}}=(X_{1},X_{2},...,X_{k})$ and $A_{2}$ as $X^{A_{2}}=(X^{\prime}_{1},X^{\prime}_{2},...,X^{\prime}_{k})$, with $X_{1}$ being the treatment variable. In the control group, the realized value of $X_{1}$ is $x^{c}_{1}$, and in the treatment group, it is $x^{t}_{1}$.
In this experimental setup, individuals in the control group exhibit a certain utility difference between $A_{1}$ and $A_{2}$, which can be represented as:
$V_{c}^{\text{exp}}$ $=V_{c}^{\text{exp}}(A_{1})-V_{c}^{\text{exp}}(A_{2})$
$=\alpha^{\prime}_{1}u_{1}(x^{c}_{1},\theta^{1}_{i})+\alpha^{\prime}_{2}u_{2}(x_{2},\theta^{2}_{i})+\ldots+\alpha^{\prime}_{k}u_{k}(x_{k},\theta^{k}_{i})-V_{c}^{\text{exp}}(A_{2}),$
where $\alpha^{\prime}_{1}+...+\alpha^{\prime}_{k}=1$. The superscript $exp$ denotes the experimental condition, and the subscript $c$ denotes the control group. Similarly, for the same individual in the treatment group, the utility evaluation can be represented as:
$V_{t}^{\text{exp}}=\alpha^{\prime}_{1}u_{1}(x^{t}_{1})+\alpha^{\prime}_{2}u_{2}(x_{2})+\ldots+\alpha^{\prime}_{k}u_{k}(x_{k})-V_{t}^{\text{exp}}(A_{2}),$
where $\sum_{j=1}^{k}\alpha^{\prime}_{j}=1$ (We omit $\theta$ because it will not affect the analysis.). In this discussion, we neglect the salience effect; thus, the salience $\alpha^{\prime}$ remain constant, implying $V_{c}^{\text{exp}}(A_{2})=V_{t}^{\text{exp}}(A_{2})$.
Therefore, the individual-level causal effect in the experiment is defined as:
$ITE^{\text{exp}}=V^{\text{exp}}(x^{t}_{1})-V^{\text{exp}}(x^{c}_{1})=\alpha^{\prime}_{1}[u_{1}(x^{t}_{1})-u_{1}(x^{c}_{1})].$
Given the assumption of no salience effect, $V(A_{2})$ cancels out, illustrating that the ITE remains the same regardless of whether the DM is comparing two alternatives in conjoint experiments or focusing on a single $A_{1}$ in other survey experiments.
In the real-world context, where the consideration set includes additional attributes $\{X_{k+1},\ldots,X_{m}\}$, the utility differences in the control and treatment scenarios are expressed as:
$V_{c}^{\text{real}}=\alpha_{1}u_{1}(x^{c}_{1})+\alpha_{2}u_{2}(x_{2})+\ldots+\alpha_{m}u_{m}(x_{m})-V_{c}^{\text{exp}}(A_{2})$
and
$V_{t}^{\text{real}}=\alpha_{1}u_{1}(x^{t}_{1})+\alpha_{2}u_{2}(x_{2})+\ldots+\alpha_{m}u_{m}(x_{m})-V_{t}^{\text{exp}}(A_{2}),$
where $\sum_{j=1}^{m}\alpha_{j}=1$.
Thus, in the real-world scenario, the individual-level causal effect is:
$ITE^{\text{real}}=\alpha_{1}[u_{1}(x^{t}_{1})-u_{1}(x^{c}_{1})].$
To compare the two causal effects without a salience effect, we establish a benchmark assumption that the salience ratio for any given dimension remains constant.
Assumption 2.
Consider two decision situations in which an individual has a salience vector $\alpha=(\alpha_{1},\alpha_{2},\ldots,\alpha_{m})$ and $\alpha^{\prime}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})$. For any non-zero $\alpha_{j}$, $\alpha_{k}$ and $\alpha^{\prime}_{j}$, $\alpha^{\prime}_{k}$ where $j,k\leq\min(m,n)$, the ratio $\frac{\alpha_{j}}{\alpha_{k}}=\frac{\alpha^{\prime}_{j}}{\alpha^{\prime}_{k}}$ holds.
This assumption reflects the natural idea that if, in one circumstance, a DM considers attribute $j$ (e.g., gender) to be more salient than attribute $k$ (e.g., age), that is, $\alpha_{j}>\alpha_{k}$, then in another similar circumstance, the same DM should still perceive gender as more salient than age, $\alpha^{\prime}_{j}>\alpha^{\prime}_{k}$, even if the exact values of salience could differ ($\alpha_{j}\neq\alpha^{\prime}_{j}$ and $\alpha_{k}\neq\alpha^{\prime}_{k}$). The assumption implies that their relative magnitude should remain the same. This assumption is not related to the consistency of relative importance because the latter is defined on the causal effect rather than salience.
Under this assumption, we can derive the relationship between $\alpha^{\prime}_{1}$ and $\alpha$, leading to the following result.
Proposition 1 (Amplification Bias).
Suppose assumptions 1 and 2 hold. If $u_{1}(x_{1}^{t})\neq u_{1}(x_{1}^{c})$, then the Individual (and Average) Treatment Effect in the experiment, compared to the real-world effect, is amplified by $\delta=\frac{1}{\alpha_{1}+\alpha_{2}+\ldots+\alpha_{k}}>1$.
Proof.
Proofs are provided in the SI. ∎
Generally, the experimental effect is larger than the real-world effect. The amplification factor $\delta$ depends on the total salience of the attributes (measured by the real-world salience $\alpha$) included in the experiment. If DMs actually care more about attributes not included, then the experimental causal effect is significantly amplified. We summarize this in the following corollary.
Corollary 1.
The Amplification Bias in the experiment, $\delta$, increases with the salience of attributes (measured by the real-world salience) excluded from the experiment.
This result is intuitive. In the experiment, DMs disproportionately focus on the attributes provided by researchers, amplifying the causal effect of each attribute compared to the real-world causal effect, where the effect is diluted by other attributes. Because the real-world consideration set is not directly observable, we cannot test this theory directly. However, if our theory holds, we can derive the following testable hypothesis from the above corollary.
Hypothesis 1.
The experimental causal effect decreases as more attributes are included.
We will test this hypothesis in the following subsection. Specifically, following common practice, the experimental causal effect is measured by the AMCE. As noted in the SI 4, since AMCE is essentially a weighted sum of ITE, the amplification bias of ITE is also reflected in AMCE.
2.1.1 Testing the Bias Amplification Hypothesis
2.1.2 Meta Analysis
To empirically test our hypotheses, we utilize a dataset comprising 67 candidate choice conjoint and vignette experiments aggregated by Schwarz and Coppock (2022). These experiments share two core characteristics: (1) a gender attribute is randomly assigned across all experiments, and (2) the outcome variable is binary, or can be transformed into a binary choice for the candidate. This design allows us to isolate and analyze the average causal effect of the gender attribute. Additionally, we gathered data on the number of non-gender attributes presented in each experiment, noting that the minimum number of attributes provided in some experiments was as low as two.
To investigate the hypothesized relationship between the number of attributes and the AMCE of gender across these experiments, we conducted a meta-analysis. Figure 2 presents the meta-regression results, employing a random effects model, for all countries and specifically for the USA—given that a majority of the experiments were conducted in the United States. The vertical line in the figure denotes the absolute value of the estimated effect, acknowledging that some experimental causal effects of gender may manifest as negative.  [^9] Consistent with our theoretical framework, a true positive effect is expected to diminish with an increasing number of attributes, whereas a true negative effect would exhibit an increasing magnitude (because its absolute value decreases. In both analytical scenarios, there is a significant negative correlation between the number of attributes and the experimental causal effect of gender, corroborating our theoretical predictions. Excluding experiments with a small number of negative estimates and outliers does not alter the results.
Figure 2: Test hypothesis of amplified effect by meta-regression. Each point represents the estimated gender effect from the conjoint experiment conducted within different countries. The size of the point denotes the inverse of the variance. The fitted meta-regression line shows that the treatment effect is decreasing with the number of attributes. Full Results are shown in the table A.1, Column 1-2.
2.1.3 Candidate-Choice Experiment
In order to address the inherent variability across different experiments and to directly test our hypothesis regarding the influence of attribute number on experimental effects, we designed and conducted a separate candidate-choice experiment maintaining a consistent set of attributes. The experiment was fielded to 1,200 US respondents by Lucid in September 2023. Lucid employs quota sampling strategy to align its participant demographics with those of the U.S. Census. Previous research by Coppock and McClellan (2019) suggests that participants recruited by Lucid exhibit behaviors comparable to those observed in respondents from representative samples across various experimental benchmark surveys. Further details about the experimental design are provided in SI D. Each respondent was randomly allocated to one of five groups. In each group, participants engaged in six paired candidate choice experiments. These groups varied in the number of attributes presented, but the number of attributes are fixed in each group.
- Group 1 was presented with only two attributes: gender and age. - Group 2 included the previous attributes plus education and tax policy, totaling four attributes. - Group 3 added race and income to the attributes in Group 2, resulting in six attributes. - Group 4 included military service and religious beliefs, bringing the total to eight attributes. - Group 5 encompassed ten attributes by adding children and marital status to those in Group 4.
Aligning with our meta-analytical focus, we specifically examined the gender effect across these groups. The results, depicted in Figure 3, in spite of 5 observations, reveal a statistically significant negative trend between the number of attributes and the magnitude of experimental effects on gender. This finding supports Hypothesis 1 regarding the diminishing effect size as the number of attributes increases. Interestingly, we observed that the gender effect in Group 4 (with eight attributes) was anomalously larger than anticipated. While this could be attributed to random variation, an alternative explanation may relate to the salience effect, which we will explore in subsequent sections.
Figure 3: Test hypothesis of amplified effect with single conjoint experiment. Given different number of attributes, the point represents the estimated AMCE of the gender effect. The size of the point denotes the inverse of the variance. the The fitted line is estimated by meta-regression. Full Results are shown in the table A.1, Column 3 and in the table A.2.
3 Salience Effect
Exploring further, we investigate the distortion in the linkage between experimental causal effects and real-world implications, moving beyond the assumption of fixed salience in the previous section. We adopt a psychological framework developed by Bordalo, Gennaioli and Shleifer (2012) to elucidate how variations in attribute salience can influence this relationship. This approach posits that individuals disproportionately weight more salient attributes when evaluating alternatives.
We introduce the concept of salience through a function that captures the degree of emphasis an individual places on a particular attribute relative to a reference point. Consider candidates $A$ and $B$, with respective attribute profiles $X^{A}=(x_{1},x_{2})$ and $X^{B}=(x^{\prime}_{1},x^{\prime}_{2})$, where we might imagine $X_{1}$ represents gender and $X_{2}$ denotes education level. The salience of an attribute hinges on its contrast with a reference, which we define as the average attribute levels, $(\overline{X}_{1},\overline{X}_{2})=\left(\frac{x_{1}+x^{\prime}_{1}}{2},\frac{x_{2}+x^{\prime}_{2}}{2}\right)$. Accordingly, the salience function for attribute $k$, denoted $\sigma(x_{k},\overline{X}_{k})$, is a measure of the ‘distance’ between attribute $j$ and its reference point.
This function aptly captures psychological perceptions of distance. A more formal discussion is available in Bordalo, Gennaioli and Shleifer (2012). The function follows that for any two intervals $[x,y]$ and $[x^{\prime},y^{\prime}]$, with $[x,y]$ fully contained within $[x^{\prime},y^{\prime}]$, it holds that $\sigma(x,y)<\sigma(x^{\prime},y^{\prime})$. This aligns with the psychological principle that higher contrasts are more perceptually salient. We also assume Homogeneity of Degree Zero: The function satisfies $\sigma(bx,by)=\sigma(x,y)$ for any positive scalar $b$. A viable form for $\sigma$ meeting these criteria is $\sigma(x,y)=\frac{|x-y|}{y}$.
An attribute is deemed salient for a candidate if it markedly stands out compared to other aspects. For instance, $\sigma(x_{1},\overline{X}_{1})$ quantifies the salience of the gender attribute for candidate $A$, and $\sigma(x_{2},\overline{X}_{2})$ does so for the education attribute. The attribute with higher salience, indicated by a greater $\sigma$ value, is considered more influential in the decision-making process.
Subsequently, individuals adjust their initial attribute weights $\alpha_{i}$ in the utility function $V_{i}=\sum_{k=1}^{K}\alpha_{k}u_{k}(X_{k},\theta_{i})$ based on the salience rankings, where $r_{i}$ represents the rank of attribute $k$’s salience (with lower $r_{k}$ indicating higher salience). For instance, should the education attribute be more salient than gender for candidate $A$, i.e., $\sigma(x_{1},\overline{X}_{1})<\sigma(x_{2},\overline{X}_{2})$, we would have $r_{1}=2$ and $r_{2}=1$. The salience of each attribute then is adjusted by $\delta^{r_{k}-1}$, where $\delta\in(0,1]$ is a parameter that denotes the degree of the severity of salient thinking. Suppose there are only two attributes, the utility function with adjusted salience has the following form:
$V=\begin{cases}\frac{\alpha_{1}}{\alpha_{1}+\delta\alpha_{2}}u_{1}(x_{1k},\theta_{1})+\frac{\delta\alpha_{2}}{\alpha_{1}+\delta\alpha_{2}}u_{2}(x_{2k},\theta_{2})&\text{if }\sigma(x_{1k},\overline{x}_{1})>\sigma(x_{2k},\overline{x}_{2})\\ \frac{\delta\alpha_{1}}{\delta\alpha_{1}+\alpha_{2}}u_{1}(x_{1k},\theta_{1})+\frac{\alpha_{2}}{\delta\alpha_{1}+\alpha_{2}}u_{2}(x_{2k},\theta_{2})&\text{if }\sigma(x_{1k},\overline{x}_{1})<\sigma(x_{2k},\overline{x}_{2})\\ \alpha_{1}u_{1}(x_{1k},\theta_{1})+\alpha_{2}u_{2}(x_{2k},\theta_{2})&\text{if}\sigma(x_{1k},\overline{x}_{1})=\sigma(x_{2k},\overline{x}_{2})\end{cases}$ (2)
Accordingly, when an attribute, such as $x_{1}$, is deemed salient, its weight in the utility function, $\alpha_{1}$, is effectively increased to $\frac{\alpha_{1}}{\alpha_{1}+\delta\alpha_{2}}$, while the weight of a less salient attribute, $\alpha_{2}$, diminishes to $\frac{\delta\alpha_{2}}{\alpha_{1}+\delta\alpha_{2}}$ (The denominator makes sure that the total salience is one.). The parameter $\delta$ serves as a measure of this effect, where a value of $\delta=1$ indicates no salience effect, implying that all attributes are weighted equally irrespective of their standout features. To encapsulate the potential for dynamic changes in attribute salience as perceived by the DM, we introduce the following assumption:
Assumption 3 (Salience Effect).
$\delta<1$.
The salience model underscores the idea that the salience attributed to each feature by an individual hinges on the extent to which the realized attribute level deviates from expectations or norms, rendering it surprising or noteworthy. With the acknowledgment of salience effects, we delve into its implications, exploring how salience effect, combined with attention effect, can significantly distort the consistency between experimental and real-world effect.
3.1 Inconsistency of Effect Direction
In previous section, we show that experimental findings can be significantly distorted due to limited attention, leading to disproportionately amplified causal effects. However, such amplification can be advantageous for researchers aiming to confirm the directionality of theoretical effects since it can reduce the required sample size for achieving a certain statistical power.
Then, a conservative but more critical requirement for experimental results to meaningfully validate theoretical predictions is the consistency of effect sign: the experimental causal effect should have the same effect direction as the real-world one (Slough and Tyson, 2024). However, the presence of salience effects introduces a risk of effect sign reversal, undermining the reliability of experimental outcomes as indicators of theoretical truth.
As before, let us still consider the ITE in an experimental context with $k$ attributes and the treatment attribute is $X_{1}$. It is evident from equation 2 that salience is influenced by the levels of attributes. Therefore, the salience rank for attribute $k$ in the $j^{th}$ alternative, $r^{j}_{k}$, is treated as a function of the treatment, $r^{j}_{k}(X_{1})$. The following proposition specifies the sufficient condition for effect sign reversal.
Proposition 2 (Effect Sign Reversal).
Suppose assumptions 2 and 3 hold. Experimental individual and average treatment effect directions are different from the real-world effects if $\exists j$ and $k$, such that $r^{j}_{k}(x_{1})\neq r^{j}_{k}(\tilde{x}_{1})$.
The critical condition within this proposition, “$\exists j$ and $k$, such that $r^{j}_{k}(x_{1})\neq r^{j}_{k}(\tilde{x}_{1})$,” elucidates a scenario where altering the treatment level affects the salience of at least one attribute within one profile. In the accompanying proof (detailed further in the SI 2), it is demonstrated that such a reversal can occur irrespective of the utility magnitudes associated with excluded attributes, or the inherent salience of the treatment attribute ($X_{1}$). This finding has profound implications for experimental design and interpretation, particularly in contexts where the primary interest lies in understanding the direction of causal effects. It warns researchers of the possibility that, under the influence of salience effects, the observed direction may actually invert.
To gain intuition regarding the phenomenon of Effect Sign Reversal, consider two hypothetical worlds, as depicted in the left panel of Table 2. In Experiment 1, we compare $B_{1}$ and $B_{2}$, each with three attributes. In Experiment 2, the comparison involves two-attribute profiles $A_{1}$ and $A_{2}$. We are particularly interested in the treatment effect of attribute $X_{1}$ as it changes from $x_{1}$ to $\tilde{x}_{1}$. Consequently, $B_{2}$ (and $A_{2}$), serving as the controlled profiles, are fixed at $X^{c}=(x^{\prime}_{1},x^{\prime}_{2},x^{\prime}_{3})$ (and ($x^{\prime}_{1},x^{\prime}_{2}$)) respectively.
We use $\beta_{k}$ ($\alpha_{k}$) to denote the salience of each attribute $x_{k}$ when the DM evaluates these alternatives. The corresponding sets of salience are illustrated in the right panel of Table 2. Assume the existence of prior salience for each attribute, denoted by $\beta=(\beta^{0}_{1},\beta^{0}_{2},\beta^{0}_{3})$ and $\alpha=(\alpha^{0}_{1},\alpha^{0}_{2})$. Consider how salience is formed and evolves as the DM observes the realized attributes during comparison. When an individual compares two profiles, differing levels of each attribute will “distort” the original salience based on the rule previously mentioned. Specifically, in World 1, when comparing $B_{1}=(x_{1},x_{2},x_{3})$ to $B_{2}=(x^{\prime}_{1},x^{\prime}_{2},x^{\prime}_{3})$, for each attribute $j$, we compute the salience function $\sigma_{k}(\cdot,\frac{x_{k}+x^{\prime}_{k}}{2})$. The original salience $\beta^{0}$ is then discounted by $\delta^{r_{k}-1}$, where $r_{k}$ represents the relative salience rank introduced earlier. We assume that the updated salience attached to $B_{1}$ follows $\beta_{1}>\beta_{2}>\beta_{3}$ and the salience attached to $B_{2}$ follows $\beta^{\prime}_{1}>\beta^{\prime}_{2}>\beta^{\prime}_{3}$.
In the hypothetical World 2, only the level of attribute 1 in the treatment group is altered, from $x_{1}$ to $\tilde{x}_{1}$. This modification affects the reference level for attribute $X_{1}$ and, consequently, the value of the salience function $\sigma_{1}$ for attribute 1. For instance, the updated salience attached to $\tilde{B}_{1}$ is now $\tilde{\beta}_{2}>\tilde{\beta}_{1}>\tilde{\beta}_{3}$; in other words, the salience of attribute 1 is now less than that of attribute 2. For simplicity, we assume that the salience values for the controlled profile remain the same as in the first comparison.
Experiment 1
World 1 World 2
$B_{1}$ $B_{2}$ $\tilde{B}_{1}$ $B_{2}$
$x_{1}$ $x^{\prime}_{1}$ $\tilde{x}_{1}$ $x^{\prime}_{1}$
$x_{2}$ $x^{\prime}_{2}$ $x^{t}_{2}$ $x^{\prime}_{2}$
$x_{3}$ $x^{\prime}_{3}$ $x^{t}_{3}$ $x^{\prime}_{3}$
Experiment 1
World 1 World 2
$\beta_{1}$ $\beta_{2}$ $\tilde{\beta_{1}}$ $\beta_{2}$
$\beta_{1}$ $\beta^{\prime}_{1}$ $\tilde{\beta}_{1}$ $\beta^{\prime}_{1}$
$\beta_{2}$ $\beta^{\prime}_{2}$ $\tilde{\beta}_{2}$ $\beta^{\prime}_{2}$
$\beta_{3}$ $\beta^{\prime}_{3}$ $\tilde{\beta}_{3}$ $\beta^{\prime}_{3}$
Experiment 2
World 1 World 2
$A_{1}$ $A_{2}$ $\tilde{A}_{1}$ $A_{2}$
$x_{1}$ $x^{\prime}_{1}$ $\tilde{x}_{1}$ $x^{\prime}_{1}$
$x_{2}$ $x^{\prime}_{2}$ $x^{t}_{2}$ $x^{\prime}_{2}$
Experiment 2
World 1 World 2
$\alpha_{1}$ $\alpha_{2}$ $\tilde{\alpha_{1}}$ $\alpha_{2}$
$\alpha_{1}$ $\alpha^{\prime}_{1}$ $\tilde{\alpha}_{1}$ $\alpha^{\prime}_{1}$
$\alpha_{2}$ $\alpha^{\prime}_{2}$ $\tilde{\alpha}_{2}$ $\alpha^{\prime}_{2}$
Table 2: Illustration of Salience Effect. $A_{j}$ and $B_{j}$ are alternatives, $x_{k}$ denotes the attribute, and $\alpha_{k}$ and $\beta_{k}$ denotes thesalience.
For Experiment 1, the utility differences in hypothetical World 1 are given by $\sum_{k}\beta_{k}u_{k}(x_{k})-\sum_{k}\beta^{\prime}_{k}u_{k}(x^{\prime}_{k})$, and in hypothetical World 2, they are $\sum_{j}\tilde{\beta}_{k}u_{k}(\tilde{x}_{k})-\sum_{k}\beta^{\prime}_{k}u_{k}(x^{\prime}_{k})$, where $\tilde{x}_{2}=x_{2}$ and $\tilde{x}_{3}=x_{3}$. Consequently, the individual treatment effect of attribute $X_{1}$ can be expressed as:
$ITE_{3}$ $=\sum_{k=1}^{3}[\beta_{k}u_{k}(x_{k})-\tilde{\beta}_{k}u_{k}(\tilde{x}_{k})]$
$=\sum_{k=1}^{2}[\beta_{k}u_{k}(x_{k})-\tilde{\beta}_{k}u_{k}(\tilde{x}_{k})]+[\beta_{3}u_{3}(x_{3})-\tilde{\beta}_{3}u_{3}(\tilde{x}_{3})]$
where the subscript $3$ indicates that a total of three attributes are considered in Experiment 1. Similarly, the individual treatment effect for Experiment 2 is
$ITE_{2}=\sum_{k=1}^{2}[\alpha_{k}u_{k}(x_{k})-\tilde{\alpha}_{k}u_{k}(\tilde{x}_{k})]$
It is evident that $ITE_{3}$ includes an additional term $[\beta_{3}u_{3}(x_{3})-\tilde{\beta}_{3}u_{3}(\tilde{x}_{3})]$ compared to $ITE_{2}$. Furthermore, the salience values denoted by $\beta$ in $ITE_{3}$ differ from those denoted by $\alpha$ in $ITE_{2}$. These two factors could lead to $ITE_{3}$ and $ITE_{2}$ having different signs.
Based on Proposition 2, we formulate the following testable hypothesis:
Hypothesis 2.
The direction of attribute effects reverses with the different numbers of attributes in a conjoint experiment.
We do not assert that an effect sign reversal will inevitably occur. For instance, if one attribute—or its salience—predominantly overshadows others in the evaluation process, the reversal of this attribute’s effect is unlikely to be observed. The reasoning behind this is straightforward: if a particular attribute (consider $X_{1}$) is significantly more important than any other attributes, it will be assigned a very large weight $\alpha_{1}$, and, typically, it will also possess a large utility value $u(X_{1})$. Therefore, in the evaluation $V$, even if the order of salience changes, the product $\alpha_{1}u(X_{1})$ remains considerably large, overshadowing other terms in the evaluation. The direct consequence is that its sign will predominantly determine the sign of $V$, and thus the ITE.
Another scenario that could preclude the observation of an effect sign reversal is the non-fulfillment of the rank change assumption. As highlighted by Proposition 2, the condition of salience rank change is critical for the observed result. The proposition below clarifies that without this condition, effect sign reversal cannot be observed under any circumstances.
Proposition 3.
Assuming conditions 2 and 3 are satisfied, if the relative salience rank $r^{j}_{k}(x_{1})=r^{j}_{k}(\tilde{x}_{1})$ for all $j$ and $k$, then the effect sign reversal will not occur.
When can we expect the salience order to remain unchanged, as presumed in the aforementioned proposition? The mechanism behind salience effects suggests that the rank of salience is contingent upon the value of the salience function, $\sigma(x_{k},\overline{X}_{k})$. Should the treatment attribute $X_{1}$ undergo a level change that is not substantial enough to significantly alter $\overline{X}_{k}$, then $\sigma(x_{k},\overline{X}_{k})$ will approximate $\sigma(\tilde{x}_{k},\overline{X}_{k})$. Consequently, the variation in salience will be insufficient to modify its relative rank. Drawing on this reasoning, the following hypothesis is proposed:
Hypothesis 3.
Attribute effect direction reversal is less likely if the level of attributes changes only marginally.
3.1.1 Testing the Effect Reversal Hypotheses
We utilize data from a conjoint experiment on hotel rooms conducted by Bansak et al. (2021) to test Hypothesis 2 . They identified four core attributes of hotel rooms as follows: “view from the room (ocean or mountain view), floor (top, club lounge, or gym and spa floor), bedroom furniture (1 king bed and 1 small couch or 1 queen bed and 1 large couch), and the type of in-room wireless internet (free standard or paid high-bandwidth wireless).” In addition to these core attributes, 18 supplementary attributes, unrelated to the core four, were included in the study. Respondents were required to select their preferred hotel room from 15 pairs of profiles that varied in attributes, each encompassing the four core attributes plus a randomly selected set of other attributes. Consequently, respondents were randomly assigned to one of 11 different experimental groups, which included profiles with 4, 5, 6, 7, 8, 9, 10, 12, 14, 18, or up to 22 attributes.
Figure 4 presents a heatmap of the results (Figure A.3 in SI labels statistical significance). The horizontal axis represents the number of attributes, and the vertical axis denotes the level of each attribute. Dark colors indicate negative AMCEs, while light colors signify positive AMCEs. The results corroborate our hypothesis on effect sign reversal (Hypothesis 2). For instance, the effects of certain variables, such as the menu, bar, closet, or pillow, are notably unstable. However, others, including View, Towels, and Internet, demonstrate considerable stability. This observation aligns with our theoretical framework, which posits that these variables are among the most critical factors in decision-making evaluations of a hotel, possessing sufficiently high prior salience and utility to dominate over other attributes. This evidence further solidifies the notion that View and Internet are indeed core attributes, as described by Bansak et al. (2021).
To test Hypothesis 3, we adapted our experimental design for candidate choice. To control for salience, instead of randomly assigning treatment levels to two profiles, we ensured that the level differences for each attribute were minimal. For instance, with age comprising five levels (40, 52, 60, 68, 75), only adjacent values could appear in a comparison. If 52 was assigned to one profile, only 40 or 60 could be assigned to the other.
Similar to the previous experiment, this one was fielded to 1,200 US respondents by Lucid in November 2023, with five pairs to evaluate this time. Table 3 displays the AMCE for gender in both the original and modified experiments. When attribute levels were assigned randomly—without controlling for salience effects—the corresponding AMCEs, as shown in the second column, experienced a sign reversal after the number of attributes increased from six to eight. The third column illustrates the effects when we controlled for salience as described. A key observation is the absence of a clear effect reversal in the reduced-salience setting. We note that this result does not suggest that researchers should assign adjacent levels for all conjoint experiments. The purpose of this design choice is to test the hypotheses derived from the theory. The broader implication on research design is discussed in the section 4.
Num Salience Effect Reduced Salience
2 0.15 0.45
4 0.10 0.10
6 0.03 0.15
8 -0.08 0.10
10 -0.02 0.27
Table 3: Gender Effect. Full results are in the table A.2 and A.3
Figure 4: Testing Hypothesis of Effect Reversal. The horizontal line denotes the number of attributes and each row illustrates the effect sign of each attribute. The dark color represents negative sign and light color represents the positive sign. The total 385 estimates are from the same model in Bansak et al. (2021)
3.2 Inconsistency of Attribute Importance
One of the principal advantages of conjoint experiments is their ability to examine the causal effects of multiple treatments simultaneously. This methodology enables researchers to assess the importance of each attribute comprehensively. The determination of relative importance is not only crucial for theoretical development but also has practical implications for policy formulation. For instance, understanding the relative importance of attributes in a candidate choice experiment can provide real-world candidates with insights into effective strategies for election campaigns. However, this application necessitates that the significance of attributes within the experimental context aligns closely with their effects in real-world scenarios.
A key insight from the discussion on salience effects is their potential to create correlations between attributes that are theoretically independent. Such correlations can distort the causal effect of an attribute, leading to a reversal of the effect sign. In a similar vein, it is reasonable to anticipate that the correlation induced by salience could also alter the perceived importance of attributes.
Proposition 4 (Attribute Importance Reversal).
Suppose assumptions 2 and 3 hold. Experimental attribute importance is different from the real-world importance if $\exists j$ and $k$, such that $r^{j}_{k}(x_{1})\neq r^{j}_{k}(\tilde{x}_{1})$.
In applications, researchers assess relative importance by comparing average marginal component effects, for example, in the studies conducted by Hainmueller, Hangartner and Yamamoto (2015) and Lehrer, Stöckle and Juhl (2024). Additionally, the use of $R^{2}$ in linear regression models offers an alternative measure of importance that circumvents the need to address the sign of the effect directly. We use both methods to test the following hypothesis:
Hypothesis 4.
The perceived importance of attributes changes with different number of attributes included in a conjoint experiment.
It is important to note, however, that not all attributes are susceptible to this phenomenon, just as not all attributes will experience a reversal in effect sign. Attributes that significantly overshadow others in terms of influence are likely to maintain their relative importance consistently.
3.2.1 Testing the hypothesis
To test the hypothesis, we utilize the previously mentioned data from the hotel rooms conjoint experiments. Recall that there are 11 experiment groups, featuring 4, 5, 6, 7, 8, 9, 10, 12, 14, 18, or 22 attributes. For each group, we estimate the AMCE and corresponding p values. To facilitate a clearer visualization of the results, we assign each attribute a rank based on its absolute magnitude within each group. In the heatmap 5 (and (figure A.4 in SI), each column presents the relative importance rank of attributes in a specific experiment group, as listed on the x-axis. The colors denote the importance rank. Also, we construct importance by $R^{2}$ as Bansak et al. (2018). Figure A.5 in SI shows the similar results.
Consistent with the hypothesis, most rows (attributes) exhibit varying colors, indicating that the relative importance of most attributes changes when the number of attributes in the experiments varies. Additionally, from Section 3.1.1, we identified View, Towels, and Internet as core variables. Our findings similarly indicate that the relative importance of these variables remains relatively stable as well. Notably, the attribute of Internet maintains its position as the most important attribute throughout, which is understandable given its perceived necessity by most individuals.
Figure 5: Testing Hypothesis of Importance Reversal. The horizontal line denotes the number of attributes and each row illustrates the relative importance of each attribute. The darker color represents less importance. The total 385 estimates are from the same model in Bansak et al. (2021).
4 Discussion
We have both theoretically and empirically demonstrated that limited attention and salience effects can lead to inconsistencies between experimental effects and their real-world counterparts. These discrepancies can manifest as differences in effect magnitude, direction, and relative importance. In this section, we delve deeper into the correlation and differences between these two types of effects and offer recommendations for refining experimental design.
Firstly, it is essential to acknowledge that salience effects are not inherently detrimental to experimental design. Whether individuals are making decisions in the real world or within an experimental framework, salience effects invariably play a role in multi-dimensional decision-making processes. As such, the objective should not be to entirely eliminate salience effects from experiments. Rather, our focus should be on ensuring that the salience effects present in the experiment closely mirrors those encountered in real-world scenarios. It is the artificially induced salience effects that pose problems. Note that in the absence of attention effects—implying that the experiment perfectly replicates the real-world decision-making environment—experimental ITEs would align with real-world effects due to the accurate matching of salience effects. However, the main challenge is that achieving this ideal scenario is often impractical. Furthermore, our focus on Average Effects (AMCE) rather than Individual Effects (ITE) necessitates a particular consideration. As outlined in the SI 4, AMCEs are derived by marginalizing over attribute components and individuals, typically requiring participants to make several rounds of decisions among multiple profiles in a conjoint experiment, with each round featuring randomly generated profiles. If the randomly assigned attribute levels are uniquely distinct and lack counterparts in the real world, this can lead to distorted salience effects. This observation leads to our first recommendation for experimental design.
Recommendation 1: Ensure that the attribute levels and their combinations within profiles accurately reflect real-world scenarios.
Secondly, in our model, attention effects are a primary cause of the inconsistencies observed in experimental outcomes. Our propositions are fundamentally anchored in the assumption of limited attention, which posits that the attributes included in an experiment may not align with those deemed relevant by DMs in real-world settings. An immediate implication for research design is that researchers should endeavor to identify which attributes DMs truly value in real-life situations, thereby ensuring a match between the consideration sets in experiment and real-world contexts. As shown by Hainmueller, Hangartner and Yamamoto (2014), if the consideration sets are matched, survey experiments have ecological validity. This alignment requires a meticulous preliminary survey prior to the experimental design. It is important to note that this does not advocate for the inclusion of an excessive number of attributes, especially those that are superfluous. While the presence of an unnecessary attribute may be harmless if it has no impact on the evaluation function, even a minimal effect from such an attribute can significantly distort outcomes when coupled with salience effects. For example, when a DM encounters an unexpected attribute in a conjoint experiment, this surprise can lead the DM to assign disproportionately high salience to this attribute, despite its negligible importance in a real-world context. This observation leads us to our second recommendation for experimental design:
Recommendation 2: Prior to designing an experiment, researchers should identify and include only the necessary set of attributes. The inclusion of irrelevant attributes can be detrimental.
The interplay between attention and salience effects raises concerns regarding the reliability of average treatment effects in experiments that involve multidimensional decision processes. Nevertheless, it is possible to derive valuable insights by adjusting our expectations to accommodate somewhat weaker conclusions. While a minor deviation from assumption 1 may inevitably lead to inconsistencies in effect magnitude, our primary interest often lies in understanding relative information, such as effect sign or relative importance. The discussions surrounding salience effects and the empirical results from the hypotheses testing have highlighted the resilience of certain “core” attributes. These core attributes, whether theoretically or empirically justified, play a pivotal role in decision-making, exerting a dominant influence over others. Consequently, the effect sign and relative importance associated with these attributes tend to remain stable.
Recommendation 3: Conjoint experiments should prioritize the investigation of core attributes. This ensures the consistency of crucial metrics like effect sign and relative importance, even amidst the complexities introduced by attention and salience effects.
Implementing the recommendations above requires researchers to adopt a phased approach of research design that underscores realism, relevance, and precision. The initial step in this process should involve an exhaustive examination of the existing literature and data pertaining to the decision-making context under investigation. This preliminary research should be complemented by exploratory interviews or surveys conducted with individuals who partake in similar decision-making activities in real-world settings. The insights garnered from these explorations are instrumental in identifying the attributes and their corresponding levels that are most pertinent to the decision-making process.
Once the relevant attributes and levels have been delineated, the subsequent phase would entail the creation of profiles that amalgamate these attribute levels in a manner reflective of real-world choices. These profiles should be both plausible and representative of typical scenarios encountered in the real world. To ensure the authenticity of these profiles, it is crucial to seek validation from experts or members of the target audience. Such validation serves as a litmus test for the realism of the profiles and their alignment with real-world experiences.
Pilot testing of the experiment with a select sample of participants from the target population would serve as a valuable step in this process. Feedback obtained from the pilot test could be used for fine-tuning the relevance and realism of the attributes and profiles. Based on this feedback, necessary adjustments can be made to the experimental design to ensure its close resemblance to real-world decision-making contexts.
In the final stage of the experimental design process, the feedback from the pilot test and expert validation is integrated to refine the design. The culmination of this process is an experimental design that is not only aligned with real-world decision-making contexts but also equipped to provide meaningful insights into the dynamics of real-world decision-making processes.
5 Concluding Remarks
An increasing number of studies employ experiments to establish causal effects within the social sciences (Druckman et al. 2006). Beyond the assurance of internal validity, external validity remains a longstanding concern. This article delves into a more fundamental aspect of generalizability: ecological validity. We theoretically and empirically demonstrate that experiments engaging in multi-dimensional decision-making processes, especially conjoint survey experiments, may lack ecological validity. In particular, we find that the magnitude of experimental effects can be inflated, and the direction of effects and their relative importance may diverge from those observed in real-world scenarios.
Through the lens of formal theory, this study uncovers how attention and salience effects could undermine ecological validity. The alignment between experimental conditions and real-world scenarios lays the groundwork for an in-depth exploration of external validity, involving variations in treatment, outcome measures, populations, and contexts. Developing a more comprehensive theory that integrates ecological validity with other forms of external validity represents a promising direction for future research.
Supplementary Information
[…]
Appendix A Formal Definition of Causal Targets
We follow Abramson et al. (2023) to develop the Average Marginal Component Effect (AMCE) and the corresponding assumptions.
In practice, decision-makers (DMs) complete $R$ rounds of choice tasks. In each round, they evaluate $J$ profiles, each with $K$ attributes. The treatment assigned to DM $i$ for the $j$th profile in the $r$th task is the $K$-dimensional vector $T_{ijr}$. For example, in round $r=1$, $T_{ijr}=T_{i11}=(x_{1},x_{2},x_{3})$ and $T_{i21}=(x^{\prime}_{1},x^{\prime}_{2},x^{\prime}_{3})$. $T_{i}$ denotes the set of treatments across all profiles and tasks. Thus, in a two-round ($R=2$) candidate choice experiment with $J=2$, $T_{i}=\{T_{i11},T_{i21},T_{i12},T_{i22}\}$.
The potential outcome in round $r$ for profile $j$, denoted by $Y_{ijr}(T_{i})$, is a function of $T_{i}$ rather than just $T_{ijr}$, without further assumptions.
The first assumption is analogous to STUVA. Let $T_{i(j)r}$ be the set that contains profiles in that round, $T_{i(j)r}=\{T_{i1r},...,T_{iJr}\}$.
Assumption 4 (Stability and No Carryover).
$Y_{ijk}(T_{i})=Y_{ijk^{\prime}}(T^{\prime}_{i})$
if $T_{i(j)r}=T^{\prime}_{i(j)r^{\prime}}$ for any $j,r$, and $r^{\prime}$.
This assumption implies that potential outcomes in round $r$ depend only on profiles in that round. Thus, we can express potential outcomes in terms of only the treatment profiles assigned in the relevant task, denoted by $Y_{ijr}(T_{i(j)r})$.
The next assumption allows us to further simplify the relationship between potential outcomes and the set of profiles.
Assumption 5 (No Profile Order Effects).
$Y_{ijr}(T_{i(j)r})=Y_{ij^{\prime}r}(T^{\prime}_{i(j^{\prime})r})$
if $T_{ijr}=T^{\prime}_{ij^{\prime}r}$, $T_{ij^{\prime}r}=T^{\prime}_{ijr}$, $T_{ijr}\neq T_{ij^{\prime}r}$, $T^{\prime}_{ijr}\neq T^{\prime}_{ij^{\prime}r}$ for any $i,j,j^{\prime},r$.
This means that simply shuffling the order of profiles in round $r$ does not affect potential outcomes as long as the profiles remain the same. Under this assumption, we can write the potential outcome as $Y_{ijr}(T_{ijr},T_{i[-j]r})$, where $T_{i[-j]r}$ is the unordered set of non-$j$th profiles.
To focus on a specific attribute, it is useful to further decompose $T_{ijr}$. We define $T_{ijrk}$ as the $k$-th attribute for profile $j$ in round $r$, and $T_{ijr[-k]}$ is the set of remaining attributes. Hence, the potential outcome can be expressed as $Y_{ijr}(T_{ijrk},T_{ijr[-k]},T_{i[-j]r})$. The two assumptions imply that the round $r$ and profile $j$ are irrelevant in most tasks, so we omit them in the potential outcome notation.
We can now define the Individual Component Effect (ICE) as follows:
Definition 2 (Individual Component Effect).
Suppose $T_{ijrk}=t_{k}$, $T_{ijrk^{\prime}}=t_{k^{\prime}}$, $T_{ijr[-k]}=t_{-k}$, $T_{i[-j]r}=\mathbf{t}$,
$ICE_{i}=Y_{i}(t_{k},t_{-k},\mathbf{t})-Y_{i}(t_{k^{\prime}},t_{-k},\mathbf{t})$
This effect reflects the change in an individual’s response when the $k$-th attribute changes while holding the other attributes constant.
In practice, most potential outcomes $Y$ are binary forced choices, represented by 00 or $1$. Theoretically, an individual chooses profile $j$ over $j^{\prime}$ in task $r$ because the utility for profile $j$ is higher than that for the latter. Hence, $Y$ is a function of the utility differences, as defined in the main text. Formally, given two profiles $j$ and $j^{\prime}$, the binary potential outcome $Y$ can be represented by $Y_{ij}=\mathbbm{1}[V_{ij}(t_{k},t_{-k},\mathbf{t})-V_{ij^{\prime}}(t_{k^{\prime}},t_{-k},\mathbf{t})\geq 0]$, where $\mathbbm{1}$ is the indicator function.
Researchers typically summarize the component effect across different profile combinations in experiments. Generally, we can marginalize over the distribution of $t_{-k}$ and $\mathbf{t}$.
Definition 3 (Marginal (Individual) Component Effect).
$MCE_{i}=\sum_{(t_{-k},\mathbf{t})\in T}[Y_{i}(t_{k},t_{-k},\mathbf{t})-Y_{i}(t_{k^{\prime}},t_{-k},\mathbf{t})]\times\mathbb{P}[T_{ijr[-k]}=t_{-k},T_{i[-j]r}=\mathbf{t}|T_{ijr[-j]},T_{i[-j]r}\in\tilde{T}]$
where $\tilde{T}$ is the intersection of the supports of $\mathbb{P}[T_{ijr[-k]}=t_{-k},T_{i[-j]r}=\mathbf{t}|T_{ijrk}=t_{k}]$ and $\mathbb{P}[T_{ijr[-k]}=t_{-k},T_{i[-j]r}=\mathbf{t}|T_{ijrk}=t_{k^{\prime}}]$.
Finally, the Average Marginal Component Effect (AMCE) is the expectation of MCE taken over the distribution of DMs.
Definition 4 (AMCE).
$AMCE$ $=\mathbb{E}[MCE_{i}]$
$=\sum_{(t_{-k},\mathbf{t})\in T}\mathbb{E}[Y_{i}(t_{k},t_{-k},\mathbf{t})-Y_{i}(t_{k^{\prime}},t_{-k},\mathbf{t})]\times\mathbb{P}[T_{ijr[-k]}=t_{-k},T_{i[-j]r}=\mathbf{t}|T_{ijr[-j]},T_{i[-j]r}\in\tilde{T}]$
Appendix B Diminish effect due to attention
In the context of unbalanced experimental designs, where the control group lacks exposure to treatment attributes, a nuanced issue arises in the estimation of causal effects. Gaines, Kuklinski and Quirk (2007) critically evaluates the logic underpinning survey experiments, suggesting the potential pitfalls of a control setup devoid of treatment attribute information.
Based on the discussion in section 3, the utility function for the control group is
$V_{i}^{exp}(x^{c}_{1})=\alpha^{\prime}_{2}u_{2}(x_{2},\theta_{2})+...+\alpha^{\prime}_{k}u_{k}(x_{2},\theta_{2}).$
For individuals in the treatment group, their utility is represented by:
$V_{i}^{exp}(x^{t}_{1})=\alpha^{\prime\prime}_{1}u_{1}(x_{1},\theta_{1})+\alpha^{\prime\prime}_{2}u_{2}(x_{2},\theta_{2})+...+\alpha^{\prime\prime}_{k}u_{k}(x_{2},\theta_{2})$
The individual causal effect of $X_{1}$ in the experiment is
$V_{i}^{exp}(x^{t}_{1})-V_{i}^{exp}(x^{c}_{1})=\alpha^{\prime\prime}_{1}u_{2}(x_{1},\theta_{1})+\sum_{j}(\alpha^{\prime\prime}_{j}-\alpha^{\prime}_{j})u_{j}(x_{2},\theta_{2}).$
This formulation uncovers at least two significant challenges. First, The objective is to isolate the causal effect attributed solely to the first dimension $X_{1}$ only. However, the term $\sum_{j}(\alpha^{\prime\prime}_{j}-\alpha^{\prime}_{j})u_{j}(x_{2},\theta_{2})$ introduces a bias that undermines this objective by capturing the differential impact of other dimensions, thereby diluting the intended causal effect of $X_{1}$. Second, the term $\alpha^{\prime\prime}_{1}u_{2}(x_{1},\theta_{1})$ is meant to capture the causal effect of $X_{1}$, considering the salience $\alpha^{\prime\prime}$ in the treatment group. Yet, the absence of a counterfactual in the control group complicates the interpretation of how the effect $X_{1}$, modulated by salience, contrasts with untreated scenarios.
Appendix C Proof of Proposition 1
Proof.
Recall, $ITE^{exp}=V^{exp}(x^{t}_{1})-V^{exp}(x^{c}_{1})=\alpha_{1}^{\prime}[u_{1}(x^{t}_{1})-u_{1}(x^{c}_{1})]$ and $ITE^{real}=\alpha_{1}[u_{1}(x^{t}_{1})-u_{1}(x^{c}_{1})]$.
The key in the proof is to re-write salience in the smaller set (exp group) $\alpha^{\prime}$ in terms of salience in the larger set $\alpha$ (real world.) Under stable salience assumption 2, we can write
$\alpha_{i}=(1-\sum_{j=k+1}^{m}\alpha_{j})\alpha_{i}^{\prime}\quad\forall i\in[1,2,...,k].$
Then, because of $u_{1}(x^{t}_{1})\neq u_{1}(x^{c}_{1})$, $\delta=\frac{ITE^{exp}}{ITE^{real}}=\frac{\alpha_{1}^{\prime}}{\alpha_{1}}=\frac{1}{\sum_{j=1}^{k}\alpha_{j}}$. The ATE follows because of the random assignment of the treatment and $ATE=\mathbb{E}[ITE]$.
Appendix D Design of Candidate-Choice Experiment
This study is pre-registered with the center for Open Science Framework (OSF). An anonymized version along with the questionnaire can be found at the OSF). The survey experiment received full Institutional Review Board (IRB) approval in Study ID [details removed for anonymous peer review] from the authors’ institutions.
D.1 Principles for Human Subjects Research
We recruit subjects using the online survey platform, Lucid Theorem, which manages relationships with suppliers who handle incentives to participants directly. Researchers pay Lucid a cost per (currently at \$1.50 for a 10–15 minute survey) for a completed interview (CPI) and Lucid pays suppliers who then provide a portion of those earnings to participants in the form of cash, gift cards, or loyalty reward points. Lucid Theorem uses a proportional sampling method to provide nationally representative samples, balancing participants based on age, gender, ethnicity, and region.
Each solicited respondent received a link from the company that redirected them to the actual survey and the consent information, hosted on a server maintained by Qualtrics. Upon completing the survey, respondents are immediately redirected back to the company’s website to claim their reward. All participants received compensation after completing the questionnaires.
In the beginning of the survey, we provided a consent form that laid out clear and comprehensive information about the project, including how data will be collected, used, and stored, the costs and benefits of participation, and contact information for the lead researcher and university IRB. Respondents were also informed that no identifying information would be collected, and that they were able to opt out of the research at any time. Each participant was required to read the consent form before proceeding with the survey (at which point consent was assumed to be granted).
The project did not involve any deception, nor did we intervene in any political processes. The data were obtained in compliance with all relevant regulations and ethical guidelines. The raw data is handled exclusively by the authors, following strict protocols to ensure confidentiality and data security.
D.2 Pre-registered Design
The online anonymized version along with the questionnaire can be found at the OSF).
This study explores how the causal effects derived from conjoint experiments can be inconsistent with the real-world effect due to attention and salience biases. Using a series of candidate choice conjoint experiments, we will randomly vary the numbers of attributes (attention bias) and their levels (salience bias) presented to the respondents and examine the differences in the average marginal component effects (AMCEs) of two key features–gender and age–common across all experimental groups. We expect that the AMCEs will be smaller the more number of attributes are included and larger when the levels of the attributes are more salient.
We conduct two candidate-choice experiment to test hypotheses outlined in the main text. The sample sizes for both experiments were 1,200, which is determined by our research budget.
Study Information
Hypotheses
H1: The AMCEs of a common attribute will be smaller the more number of attributes are included in the conjoint experiment
H2: The AMCEs of a common attribute will be larger when the levels of the attributes are more salient in the conjoint experiment.
Part 1:
In the firs experiment, respondents see a conjoint table with a hypothetical candidate that is described by K (=2,4,6,8,10) attributes as shown in the example table below:
Figure A.1: Caption
We randomly assign participants into one of five groups. For each attribute, levels are also randomly assigned. Each participant completed six rounds of the candidate choice experiment.
  • Group 1: each hypothetical candidate has K=2 attributes including Gender and Age; there are total of 6 rounds for each participant
  • Group 2: each hypothetical candidate has K=4 attributes including Gender, Age, Education, and Tax; there are total of 6 rounds for each participant
  • Group 3: each hypothetical candidate has K=6 attributes including Gender, Age, Education, Tax, Race, and Income; there are total of 6 rounds for each participant
  • Group 4: each hypothetical candidate has K=8 attributes including Gender, Age, Education, Tax, Race, Income, Religion, and Military service; there are total of 6 rounds for each participant
  • Group 5: each hypothetical candidate has K=10 attributes including Gender, Age, Education, Tax, Race, Income, Religion, Military service, Gay Marriage, and Children; there are total of 6 rounds for each participant
Attribute level:
Figure A.2: Attribute Table
Part 2:
In the second experiment, respondents are required to complete 5 rounds of candidate choice experiment similar to the first experiment. However, we control attribute salience in each pair. In particular, we let some attributes be randomly chosen so that their levels are close to each other. For example, for attribute age = [40,52,60,68,75], only levels (40,52) , (52,60), (60,68), and (68,75) will be randomly chosen. Those levels are age, income, same sex marriage, tax, education, and children.
  • If K=2, attributes are Gender and Age
  • If K=4, attributes are Gender, Age, Tax, and Education.
  • If K=6, attributes are Gender, Age, Tax, Education, Race, and Income.
  • If K=8, attributes are Gender, Age, Tax, Education, Race, Income, Military, and Religion.
  • If K=10, attributes are Gender, Age, Tax, Education, Race, Income, Military, Religion, Same Sex Marriage, and Children.
Analysis:
(1) We will calculate the average marginal component effect (binary and 10-point rate), and $R^{2}$ for each attribute in each K group. We also present both results with and without covariate adjustment.
(2) Check the relationship between AMCE and the number of attributes K by linear regression and meta-regression.
(3) Check the relationship between the treatment effect direction (positive, negative) and the number of attributes K by drawing a heatmap.
(4) Check the relationship between attribute importance (rank of $R^{2}$) and the number of attributes K by drawing a heatmap.
(5) Check the relationship between the treatment effect direction/attribute importance and the number of attributes K conditional on whether levels are salient or not.
Appendix E Proof of Proposition 2
Proof.
Without loss of generality, we consider the same setting as depicted in Table 2. This entails comparing the change in effect sign between scenarios with two and three attributes. For instances involving more than two attributes, we can aggregate these attributes into a single composite attribute with a composite salience (thanks to linearity), and the proof follows analogously.
In the proposition, the condition is ‘$\exists k$ such that $r^{j}_{k}(x_{1})\neq r^{j}_{k}(\tilde{x}_{1})$’. According to the salience mechanism, varying levels of $X_{1}$ influence its salience function and relative rank $r_{1}$. Hence, we can confidently assume $r^{1}_{1}(x_{1})\neq r^{1}_{1}(\tilde{x}_{1})$. Under this premise, as detailed in the main text, the Individual Treatment Effect (ITE) for two attributes (say, in the experiment) is
$ITE_{2}=[\alpha_{1}u_{1}(x_{1})-\tilde{\alpha}_{1}u_{1}(\tilde{x}_{1})]+(\alpha_{2}-\tilde{\alpha}_{2})u_{2}(x_{2})$
and the ITE for three attributes (say, in the real world) is
$ITE_{3}=[\beta_{1}u_{1}(x_{1})-\tilde{\beta}_{1}u_{1}(\tilde{x}_{1})]+(\beta_{2}-\tilde{\beta}_{2})u_{2}(x_{2})+(\beta_{3}-\tilde{\beta}_{3})u_{3}(x_{3})$
See the differences between the first term:
$(\alpha_{1}-\beta_{1})u_{1}(x_{1})+(\tilde{\beta}_{1}-\tilde{\alpha}_{1})u_{1}(\tilde{x}_{1})$ (3)
Without loss of generality, we normalize utility to be positive. By stable salience assumption, $\beta^{0}_{1}=(1-\beta^{0}_{3})\alpha_{1}^{0}$ and $\beta^{0}_{2}=(1-\beta^{0}_{3})\alpha_{2}^{0}$.
Note that $\alpha_{1}=\frac{\alpha_{1}^{0}}{\alpha_{1}^{0}+\delta\alpha^{0}_{2}}>\beta_{1}=\frac{\beta_{1}^{0}}{\beta_{1}^{0}+\delta\beta^{0}_{2}+\delta^{2}\beta^{0}_{3}}=\frac{\alpha^{0}_{1}}{\alpha^{0}_{1}+\delta\alpha^{0}_{2}+\frac{\beta^{0}_{3}}{1-\beta^{0}_{3}}\delta^{2}}$ and thus $\alpha_{1}-\beta_{1}>0$; similarly, $\tilde{\beta}_{1}-\tilde{\alpha}_{1}<0$.
Because the proposition concerns the existence of a particular scenario, it suffices to identify a single instance in which the effect sign reverses.
Case 1: Suppose $(\alpha_{1}-\beta_{1})u_{1}(x_{1})+(\tilde{\beta}_{1}-\tilde{\alpha}_{1})u_{1}(\tilde{x}_{1})=0$.
The second part in $ITE_{2}$ is negative because $\alpha_{2}=\frac{\delta\alpha_{2}^{0}}{\alpha_{1}^{0}+\delta\alpha^{0}_{2}}<\frac{\alpha^{0}_{2}}{\delta\alpha^{0}_{1}+\alpha^{0}_{2}}=\tilde{\alpha}_{2}$, recall $\delta\in(0,1)$ by assumption. Now, WLOG, let us assume $ITE_{2}<0$. Therefore, we need to find conditions such that $ITE_{3}\geq 0$.
We know the second part in $ITE_{3}$, $(\beta_{2}-\tilde{\beta}_{2})u_{2}(x_{2})$ must be negative because $\beta_{2}=\frac{\delta\beta^{0}_{2}}{\beta_{1}^{0}+\delta\beta^{0}_{2}+\delta^{2}\beta^{0}_{3}}<\tilde{\beta}_{2}=\frac{\beta^{0}_{2}}{\delta\beta^{0}_{1}+\beta^{0}_{2}+\delta^{2}\beta^{0}_{3}}$. Thus, in order to $ITE_{3}\geq 0$, we can let $\beta_{3}>\tilde{\beta}_{3}$. Because $\beta_{3}=\frac{\delta^{2}\beta^{0}_{3}}{\beta_{1}^{0}+\delta\beta^{0}_{2}+\delta^{2}\beta^{0}_{3}}$ and $\tilde{\beta}_{3}=\frac{\delta^{2}\beta^{0}_{3}}{\delta\beta^{0}_{1}+\beta^{0}_{2}+\delta^{2}\beta^{0}_{3}}$, simple algebra shows that
$\begin{cases}\beta_{3}>\tilde{\beta}_{3}&if\quad\beta^{0}_{1}<\beta^{0}_{2}\\ \beta_{3}<\tilde{\beta}_{3}&if\quad\beta^{0}_{1}>\beta^{0}_{2}\\ \end{cases}$
This means that, $ITE_{3}\geq 0$ if and only if $u_{3}(x_{3})\geq\frac{\tilde{\beta}_{2}-\beta_{2}}{\beta_{3}-\tilde{\beta}_{3}}u_{2}(x_{2})$ and $\beta^{0}_{1}>\beta^{0}_{2}$.
Here is another possible situation.
Case 2: Suppose $(\alpha_{1}-\beta_{1})u_{1}(x_{1})+(\tilde{\beta}_{1}-\tilde{\alpha}_{1})u_{1}(\tilde{x}_{1})<0$ and $ITE_{2}>0$.
We need find conditions such that $ITE_{3}\leq 0$. We notice the first term in $ITE_{3}$ is positive and the second term is negative.
Now suppose $\beta^{0}_{1}<\beta^{0}_{2}$ so that $\tilde{\beta}_{3}>\beta_{3}$. Therefore, in order to $ITE_{3}\leq 0$, $u_{3}(x_{3})$ must be close to 0.
As demonstrated in the proof, there exists considerable flexibility to achieve the desired outcome. Identifying just one scenario that aligns with our hypothesis is sufficient.
Appendix F Proof of Proposition 3
Proof.
Following the notational conventions established in the Proof of Proposition 2, let us consider comparing the effect sign in scenarios with $m$ and $n$ attributes, where $m<n$.
The differential in evaluation for the world with treatment $x_{1}$ is expressed as: $\sum_{k=1}^{n}\beta_{k}u_{k}(x_{k})-\beta^{\prime}_{k}u_{k}(x^{\prime}_{k})$, and for the world with treatment $\tilde{x}_{1}$, it is: $\sum_{k=1}^{n}\beta_{k}u_{k}(\tilde{x}_{k})-\beta^{\prime}_{k}u_{k}(x^{\prime}_{k})$, with the assumption that $\tilde{x}_{-1}=x_{-1}$.
Given that there is no change in the salience ranking, in accordance with the salience mechanism, the salience vectors $\beta$ and $\beta^{\prime}$ remain consistent across the two worlds despite varying levels of treatment for $X_{1}$.
Hence, the calculation of $ITE$ becomes straightforward,
$ITE_{m}=\beta_{1}[u_{1}(x_{1})-u_{1}(\tilde{x}_{1})]$
and
$ITE_{n}=\alpha_{1}[u_{1}(x_{1})-u_{1}(\tilde{x}_{1})]$
Given salience are positive, we conclude that the sign of $ITE_{m}$ and $ITE_{m}$ must be the same.
Appendix G Proof of Proposition 4
Proof.
In the scenario described, where we observe a shift in treatment value from $x^{t}_{1}$ to $\tilde{x}^{t}_{1}$, resulting in a decreased updated salience for the treatment profile, specifically $\tilde{\beta^{t}_{1}}<\tilde{\beta^{t}_{2}}$, while the salience remains unchanged when $x^{t}_{2}$ transitions to $\tilde{x}^{t}_{2}$, and salience for the control profile remains constant, our analysis can be simplified.
For the two-attribute case ((say, in the experiment)), the Individual Treatment Effect (ITE) for the first attribute can be articulated as follows:
$ITE^{1}_{2}=[\alpha_{1}u_{1}(x_{1})-\tilde{\alpha}_{1}u_{1}(\tilde{x}_{1})]+(\alpha_{2}-\tilde{\alpha}_{2})u_{2}(x_{2})$
and the ITE for the second attribute is
$ITE^{2}_{2}=\alpha_{2}[u_{2}(x_{2})-u_{2}(\tilde{x}_{2})]$
Similarly, we get two ITEs in the three-attributes case (say, in the real world),
$ITE^{1}_{3}=[\beta_{1}u_{1}(x_{1})-\tilde{\beta}_{1}u_{1}(\tilde{x}_{1})]+(\beta_{2}-\tilde{\beta}_{2})u_{2}(x_{2})+(\beta_{3}-\tilde{\beta}_{3})u_{3}(x_{3})$
and
$ITE^{2}_{3}=\beta_{2}[u_{2}(x_{2})-u_{2}(\tilde{x}_{2})]$
Again, we assume there is $u_{1}(\tilde{x_{1}})$ such that $(\alpha_{1}-\beta_{1})u_{1}(x_{1})+(\tilde{\beta}_{1}-\tilde{\alpha}_{1})u_{1}(\tilde{x}_{1})=0$.
Now, we consider the case that $|ITE^{1}_{2}|>|ITE^{2}_{2}|$. We hope to find conditions such that $|ITE^{1}_{3}|\leq|ITE^{2}_{3}|$.
Without loss of generality, we further assume $u_{2}(x_{2})>u_{2}(\tilde{x}_{2})$ so that both $ITE^{2}_{2}$ and $ITE^{2}_{3}$ are positive.
Recall $\beta_{2}<\tilde{\beta_{2}}$, therefore, we can see if $\beta_{3}>\tilde{\beta_{3}}$, which means $\beta^{0}_{1}<\beta^{0}_{2}$, we can find $u_{3}(x_{3})\in[\frac{(\tilde{\beta_{2}}-\beta_{2})u_{2}(x_{2})-ITE^{2}_{3}}{\beta_{3}-\tilde{\beta_{3}}},\frac{ITE^{2}_{3}+(\tilde{\beta_{2}}-\beta_{2})u_{2}(x_{2})}{\beta_{3}-\tilde{\beta_{3}}}]$ satisfies the condition.
Appendix H Tables
Table A.1: Gender Effects Meta-regression
All USA Own
(1) (2) (3)
Number --0.0029∗∗∗ --0.0037∗∗∗ --0.0148
(0.0009) (0.0037) (0.0054)
Intercept 0.0459∗∗∗ 0.0549∗∗∗ 0.1653∗∗
(0.0059) (0.0071) (0.0352)
$R^{2}$ 0.21 0.43 0
$\tau^{2}$ 0.0003 0.0003 0
Notes: $R^{2}$ denotes the amount of heterogeneity accounted for and $\tau^{2}$ is the estimated amount of residual heterogeneity. p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01.
Dependent Variable: Preference Score
(1) (2) (3) (4) (5)
GenderMale 0.154∗∗ 0.098 0.031 --0.081 --0.018
(0.063) (0.068) (0.067) (0.071) (0.069)
Age52 0.119 --0.041 0.032 0.004 --0.131
(0.099) (0.107) (0.107) (0.112) (0.108)
Age60 --0.162 --0.119 --0.194 --0.048 --0.110
(0.101) (0.107) (0.106) (0.111) (0.109)
Age68 --0.425∗∗∗ --0.331∗∗∗ --0.273∗∗ --0.145 --0.253∗∗
(0.101) (0.107) (0.107) (0.112) (0.108)
Age75 --1.340∗∗∗ --0.830∗∗∗ --0.566∗∗∗ --0.551∗∗∗ --0.394∗∗∗
(0.101) (0.107) (0.107) (0.112) (0.110)
EducationBA from Ivy League University 0.048 --0.027 --0.042 0.045
(0.083) (0.082) (0.087) (0.084)
EducationBA from State University 0.076 0.001 --0.036 0.026
(0.083) (0.081) (0.087) (0.085)
TaxOppose --0.327∗∗∗ --0.374∗∗∗ --0.209∗∗ --0.094
(0.083) (0.082) (0.087) (0.085)
TaxSupport 0.828∗∗∗ 0.401∗∗∗ 0.213∗∗ 0.327∗∗∗
(0.083) (0.081) (0.087) (0.084)
RaceBlack 0.137 0.116 0.037
(0.082) (0.087) (0.084)
(1) (2) (3) (4) (5)
RaceWhite 0.066 0.163 0.078
(0.082) (0.087) (0.084)
Income32,000 --0.080 0.009 --0.159
(0.117) (0.123) (0.119)
Income5.1 million --0.347∗∗∗ --0.085 --0.179
(0.116) (0.124) (0.120)
Income54,000 --0.030 0.089 0.075
(0.116) (0.123) (0.119)
Income65,000 --0.068 0.252∗∗ 0.007
(0.116) (0.123) (0.119)
Income92,000 --0.051 0.095 --0.030
(0.117) (0.123) (0.120)
MilitaryServed 0.209∗∗∗ 0.134
(0.071) (0.069)
ReligionMormon --0.429∗∗∗ --0.173∗∗
(0.087) (0.084)
ReligionProtestant --0.024 0.040
(0.087) (0.084)
Children1 0.073
Children2 0.159
(1) (2) (3) (4) (5)
Children3 0.138
Children4 0.158
MarriageOppose --0.481∗∗∗
MarriageSupport 0.010
Constant 6.240∗∗∗ 5.540∗∗∗ 5.813∗∗∗ 5.697∗∗∗ 5.375∗∗∗
(0.078) (0.108) (0.139) (0.161) (0.175)
Observations 6,300 6,187 5,825 5,713 6,180
R2 0.042 0.046 0.025 0.019 0.018
Adjusted R2 0.041 0.044 0.022 0.016 0.014
Residual Std. Error 2.513 (df = 6294) 2.668 (df = 6177) 2.548 (df = 5808) 2.680 (df = 5693) 2.702 (df = 6154)
F Statistic 55.115∗∗∗ (df = 5; 6294) 32.840∗∗∗ (df = 9; 6177) 9.211∗∗∗ (df = 16; 5808) 5.947∗∗∗ (df = 19; 5693) 4.481∗∗∗ (df = 25; 6154)
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
Table A.2: Full Results for Candidate-Choice Conjoint Experiments
Dependent Variable: Preference Score
(1) (2) (3) (4) (5)
genderMale 0.453∗∗∗ 0.098 0.153 0.103 0.272∗∗
(0.101) (0.099) (0.107) (0.097) (0.109)
age52 0.088 0.028 --0.398∗∗ --0.151 0.200
(0.173) (0.174) (0.184) (0.174) (0.193)
age60 --0.079 0.157 --0.505∗∗∗ --0.001 0.347
(0.173) (0.173) (0.184) (0.174) (0.193)
age68 --0.366∗∗ 0.128 --0.575∗∗∗ --0.039 0.383∗∗
(0.174) (0.173) (0.186) (0.174) (0.193)
age75 --0.855∗∗∗ --0.318 --1.019∗∗∗ --0.385 --0.102
(0.205) (0.214) (0.234) (0.212) (0.239)
eduBA from Ivy League University --0.792∗∗∗ --0.427∗∗ --0.287 0.096
(0.162) (0.179) (0.163) (0.192)
eduBA from State University --0.281∗∗ --0.048 --0.268∗∗ 0.132
(0.126) (0.137) (0.124) (0.143)
taxOpposite --0.184 --0.372∗∗∗ --0.483∗∗∗ --0.170
(0.129) (0.140) (0.125) (0.145)
taxSupport 1.034∗∗∗ 0.498∗∗∗ 0.342∗∗∗ 0.271
(0.126) (0.138) (0.125) (0.143)
raceWhite 0.056 0.095 --0.010
(0.108) (0.103) (0.116)
(1) (2) (3) (4) (5)
income32,000 --0.045 --0.338 0.021
(0.211) (0.186) (0.214)
income5.1 million --0.260 --0.774∗∗∗ 0.058
(0.212) (0.189) (0.217)
income54,000 --0.238 --0.344∗∗ 0.024
(0.167) (0.150) (0.171)
income65,000 --0.079 0.013 0.078
(0.166) (0.151) (0.170)
income92,000 --0.047 --0.250 0.172
(0.165) (0.151) (0.171)
miliServed 0.088 0.069
(0.101) (0.114)
reliProtestant 0.116 --0.041
(0.110) (0.126)
child1 0.320
child2 0.044
(1) (2) (3) (4) (5)
child3 0.158
child4 --0.285
gayOpposite --0.453∗∗∗
gaySupport 0.599∗∗∗
Constant 5.707∗∗∗ 5.664∗∗∗ 6.406∗∗∗ 6.071∗∗∗ 4.829∗∗∗
(0.154) (0.180) (0.222) (0.234) (0.300)
Observations 2,373 2,359 2,420 2,775 2,520
R2 0.024 0.086 0.052 0.045 0.046
Adjusted R2 0.022 0.082 0.046 0.039 0.037
Residual Std. Error 2.456 (df = 2367) 2.353 (df = 2349) 2.599 (df = 2404) 2.493 (df = 2757) 2.687 (df = 2496)
F Statistic 11.462∗∗∗ (df = 5; 2367) 24.485∗∗∗ (df = 9; 2349) 8.762∗∗∗ (df = 15; 2404) 7.611∗∗∗ (df = 17; 2757) 5.228∗∗∗ (df = 23; 2496)
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
Table A.3: Full Results for Candidate-Choice Conjoint Experiments (Control Salience)
Appendix I Figures
Figure A.3: Testing Hypothesis of Effect Reversal with P-value $<0.1$
Figure A.4: Testing Hypothesis of Importance Reversal with P-value $<0.1$
Figure A.5: Testing Hypothesis of Importance Reversal with $R^{2}$

### v2

@@
 Abstract
-Can the causal effects estimated in experiment be generalized to real-world scenarios?
+Can the causal effects estimated in an experiment be generalized to real-world scenarios?
 This question lies at the heart of social science studies.
-External validity primarily assesses whether experimental effects persist across different settings, implicitly presuming the experiment’s ecological validity—that is, the consistency of experimental effects with their real-life counterparts.
+External validity primarily assesses whether experimental effects persist across different settings, implicitly presuming the consistency of experimental effects with their real-life counterparts.
 However, we argue that this presumed consistency may not always hold, especially in experiments involving multi-dimensional decision processes, such as conjoint experiments.
@@
 Introduction
-The randomized experiment is often hailed as the gold standard for identifying and estimating causal effects within the social sciences.
-One popular such method is conjoint experiment, which is frequently employed by researchers to examine how individual preferences for candidates or immigrants are influenced by characteristics such as gender, race, experience, and skills (Eggers, Vivyan and Wagner, 2018; Carnes and Lupu, 2016; Sanbonmatsu, 2002).
+The randomized experiment is considered as the gold standard for identifying and estimating causal effects within the social sciences.
+One popular such method is the conjoint experiment, which is frequently employed by researchers to examine how individual preferences for candidates or immigrants are influenced by characteristics such as gender, race, experience, and skills (Eggers, Vivyan and Wagner, 2018; Carnes and Lupu, 2016; Sanbonmatsu, 2002).
 However, the interest of researchers extends well beyond the mere observation of causal effects within an experimental setting.
-Especially, we ask whether these experimental estimates accurately mirror real-world effects.
+Specifically, we ask whether these experimental estimates accurately mirror real-world effects.
 For instance, in a conjoint survey experiment that reveals a preference for female candidates with an average causal effect of approximately 2%, researchers hope that this finding corresponds to an actual increase in preference for female candidates in real-world scenarios.
 When such a correspondence is established, we regard the experimental effect as being consistent with the real-world effect.
-This consistency, which establishes the ecological validity of experiment, serves as a crucial concern of social science studies.
-Ecological validity represents a critical aspect of external validity, distinct from internal validity, which pertains to the validity of conclusions within the experimental setting (Kihlstrom 2021).
-As depicted in Figure 1, external validity primarily concerns the generalization of results from the experimental domain to real-world contexts with varied populations, treatments, outcomes, and contexts (Cook, Campbell and Shadish 2002).
-A fundamental premise is that the effects observed in an experiment should be consistent with real-life effects.
-If the effects estimated in the experiment do not align with those experienced by the same individuals in the real world, it undermines the rationale for discussing the generalizability of these results to other populations, treatments, outcomes, or contexts.
-Figure 1: Experimental Validity: Ecological validity as a premise or a special subset of External Validity
-We demonstrate that experiments involving multi-dimensional decision-making processes, such as the widely used vignette and conjoint survey experiments with multiple attributes manipulated at random, frequently fail to achieve ecological validity.
+We demonstrate that experiments involving multi-dimensional decision-making processes, such as the widely used vignette and conjoint survey experiments with multiple attributes manipulated at random, often fail to achieve such consistency.
 This failure raises substantial concerns about our ability to generalize experimental insights to other real-life scenarios effectively.
@@
 The Consistency of Effect Direction posits that the sign (positive or negative) of an experimental effect should match its real-world counterpart.
-Achieving magnitude consistency ensures direction consistency by default, rendering the latter a less rigorous yet still crucial requirement for real-world insights.
 Lastly, Consistency of Attribute Importance emphasizes the need for experimental outcomes to truthfully represent the relative magnitude of different attributes.
-This aspect is particularly valuable for researchers aiming to identify which attributes most significantly influence variations in outcomes, marking it as another derivative of magnitude consistency.
+This aspect is particularly valuable for researchers aiming to identify which attributes most significantly influence variations in outcomes.
 It is not imperative for an experiment to meet all three consistency criteria simultaneously; rather, the requisite consistency should align with the experiment’s objectives.
@@
 Similarly, attribute importance consistency is sufficient when the goal is to understand attribute influences and to predict outcomes [^3] In machine learning exercises, for example, predictors that account for the majority of variation are assigned greater importance and have a more significant impact on predictions..
-We develop a formal theory to elucidate why experiments involving multi-dimensional decision-making frequently fall short of meeting the three consistency criteria.
+We develop a formal theory to elucidate why and under what conditions experiments involving multi-dimensional decision-making frequently fall short of meeting the three consistency criteria.
 At the heart of our theoretical framework lies the interplay between attention and salience.
@@
 For example, a key hypothesis is that the observed effect diminishes or even reverses as the number of attributes increases.
-To empirically test these hypotheses, we analyze data from multiple existing studies and conduct additional, pre-registered conjoint experiments  [^4] The study is pre-registered and approved by authors’ Institutional Review Board; More details are in the SI D..
-The experimental findings consistently support our hypotheses, thereby reinforcing our theoretical framework.
+To empirically test these hypotheses, we analyze data from multiple existing studies and conduct additional, pre-registered conjoint experiments.
+[^4] The study is pre-registered and approved by authors’ Institutional Review Board; More details are in the SI D.
+The experimental findings consistently support our hypotheses.
 We do not claim that our proposed theory is the only reason that generates those inconsistency.
@@
 One main reason is that their experiment successfully eliminates the attention effect so that the consideration set of the experiment is almost the same to the real-world one.
-Our findings provide a theoretical explanation to understand conditions under which conjoint experiments are reliable.
+Jenke et al.
+(2021) utilize eye-tracking to examine attention dynamics and the stability of conjoint experiments.
+Their findings offer empirical support for our assumptions and theoretical framework.
+Firstly, they observe that respondents’ attention diminishes markedly as the number of attributes increases.
+Secondly, while their research concludes that conjoint experiments are substantially stable, they still note some modest variations, including effect reversals.
+Our study provides a theoretical framework to understand the conditions under which conjoint experiments are reliable or unreliable.
 Second, our work engages with the emerging literature on Theoretical Implication of Empirical Methods (TIEM) (Slough and Tyson, 2023; De Mesquita and Tyson, 2020; Slough, 2023; Fu and Slough, 2023), which critically assesses empirical methodologies through the lens of formal theory.
@@
 2015; Huang 2022).
-Our paper focuses on ecological validity, which compares the experimental effect and real-world effect even for the same individuals, treatments, and outcomes.
-Many factors can affect ecological validity, such as the famous Hawthorne effect (Adair, 1984).
-We provide a new explanation to analyze ecological validity for experiments involving multi-dimensional decision processes.
+Our paper focuses on the consistency between the experimental effect and real-world effect, even for the same individuals, treatments, and outcomes.
+Many factors can affect consistency, such as the famous Hawthorne effect (Adair, 1984).
+We provide a new explanation for experiments involving multi-dimensional decision processes.
 Lastly, we broaden the application of salience theory beyond its traditional confines in political science and economics, where its implications for election, party competition, agenda-setting, and judicial biases have been well-documented (Dragu and Fan, 2016; Riker, Riker and Riker, 1986; Ascencio and Gibilisco, 2015; Guthrie, Rachlinski and Wistrich, 2000; Viscusi, 2001).
@@
 Theoretically, the DM chooses candidate $A_{1}$ over $A_{2}$ because the DM has a higher evaluation for the former.
+Because the utility difference drives the potential outcome we observe, without loss of generality, throughout the main text, we focus on $V$ rather than $Y$  [^6] It is evident that the binary choice potential outcome can be defined as $Y_{i}^{j}=\mathbbm{1}[V^{j}_{i}\geq 0]$, where $j=1,2$.
+For outcomes that are preference scores, our results also hold because the score is also a function of the evaluation.
+In some survey experiments, the DM may only observe one alternative.
+In such cases, we can simply set $V_{i}(A_{2})=0$..
 Therefore, $Y$ is a function of the utility difference.
 The corresponding utility differences in World 1 are $V^{1}_{i}=V_{i}(A_{1})-V_{i}(A_{2})$, and in World 2, they are $V^{2}_{i}=V_{i}(\tilde{A}_{1})-V_{i}(A_{2})$  [^7] Note that $V_{i}(A_{2})$ in the two hypothetical worlds may differ due to the salience effect, which we will discuss in Section 3..
-It is evident that the binary choice potential outcome can be defined as $Y_{i}^{j}=\mathbbm{1}[V^{j}_{i}\geq 0]$, where $j=1,2$.
-Because the utility difference drives the potential outcome we observe, without loss of generality, throughout the main text, we focus on $V$ rather than $Y$  [^7] For outcomes that are preference scores, our results also hold because the score is also a function of the evaluation.
-In some survey experiments, the DM may only observe one alternative.
-In such cases, we can simply set $V_{i}(A_{2})=0$..
 Consequently, the individual causal effect for attribute $X_{1}$ when it changes from $x_{1}$ to $\tilde{x}_{1}$, given the other attributes remain constant, is defined as the difference-in-differences $V^{1}_{i}-V^{2}_{i}$.
-Definition 1 (ITE).
-Given an alternative set $A=\{A_{1},\tilde{A_{1}},A_{2}\}$ characterized by attributes $X=(X_{1},X_{2},...,X_{K})$.
-Let $X^{A_{1}}=(x_{j},x_{-j})$ represent the attributes of $A_{1}$, where $x_{j}$ is the value of attribute $X_{j}$ for $A_{1}$ and $x_{-j}$ denotes the values of all other attributes except $X_{j}$.
-Similarly, let $\tilde{X}^{\tilde{A}_{1}}=(\tilde{x}_{j},x_{-j})$ and $X^{A_{2}}=(x^{\prime}_{j},x^{\prime}_{-j})$ represent the attributes of $\tilde{A}_{1}$ and $A_{2}$.
-For a DM $i$, the individual treatment effect of attribute $X_{j}$ with respect to changing its value from $x_{j}$ to $\tilde{x}_{j}$ is defined as: $ITE^{i}=[V_{i}(X^{A_{1}})-V_{i}(X^{A_{2}})]-[V_{i}(\tilde{X}^{A_{1}})-V_{i}(X^{A_{2}})]$.
 This concept is also referred to as the individual component effect (ICE) by Abramson et al.
@@
 The formulation aligns with methodologies employed in the conjoint experiment literature, as discussed by Hainmueller, Hopkins and Yamamoto (2014).
-Given that alternatives are characterized by attributes $X$, we also use the notation $ITE^{i}=[V_{i}(A_{1})-V_{i}(A_{2})]-[V_{i}(\tilde{A}_{1})-V_{i}(A_{2})]$ to denote ITE in this paper.
-This example of ITE underscores the notion that, within a multi-dimensional decision-making environment, the ITE is contingent not only on the treatment attribute $X_{1}$ but also on other attributes.
 It is worth noting that the formal derivation of the ITE hinges on other assumptions: Stability and No Carryover Effects, and No Profile-Order Effects, which are standard in the conjoint experiment literature.
-A more formal discussion is available in the Supplementary Information (SI) 4.
-The Marginal Component Effect (MCE) is defined as a weighted sum of ITEs, and the Average Marginal Component Effect (AMCE) represents the expected value of MCE across the distribution of DMs.
+As shown by Abramson et al.
+(2023), the Marginal Component Effect (MCE) is defined as a weighted sum of ITEs over different profile combinations in experiments: $MCE_{i}=ITE_{i}\times\mathbb{P}[\text{profiles contain }x_{1}\text{and}\tilde{x}_{1}\text{occur}]$.
+Then, the Average Marginal Component Effect (AMCE) represents the expected value of MCE across the distribution of DMs: $AMCE=\mathbb{E}[MCE_{i}]$.
+A more formal discussion and derivation are available in the Supplementary Information (SI) 3.
 Our analytical framework is not confined to conjoint experiments alone; it possesses the flexibility to be applied across a spectrum of (survey) experimental designs.
@@
 2 Attention Effect
-Individuals have limited attention spans.
+Individuals have limited attention.
 Economists have observed that when people make decisions, such as purchasing a computer, they do not consider all possible options (Hausman, 2008).
-Within the literature, the subset of alternatives that a DM considers when making a decision is referred to as the “consideration set” (Hausman, 2008; Wright and Barbour, 1977).
+Within the literature, the subset of alternatives that a DM considers when making a decision is referred to as the “consideration set”, the set of attributes that DM indeed considers when making decisions (Hausman, 2008; Wright and Barbour, 1977).
 It has been well-argued and demonstrated that due to cognitive limitations, humans cannot allocate sufficient attention to every potential attribute (Stigler, 1961; Jones and Baumgartner, 2005; Chetty, Looney and Kroft, 2009).
@@
 Firstly, researchers seldom possess complete knowledge of which attributes constitute the real-world consideration set $X^{r}$.
-Secondly, even with such knowledge, it is impractical to present all relevant information about these attributes within the confines of a single survey experiment  [^8] A more general assumption could be that $X^{e}\neq X^{r}$.
-However, the empirical results in the paper provide little evidence on this assumption..
+Secondly, even with such knowledge, it is impractical to present all relevant information about these attributes within the confines of a single survey experiment.
+A more general assumption could be that $X^{e}\neq X^{r}$.
+However, the empirical results in the paper provide little evidence on this assumption.
 Without loss of generality, we assume that the real-world consideration set is $X^{r}=\{X_{1},X_{2},...,X_{m}\}$ and the experimental consideration set is $X^{e}=\{X_{1},X_{2},...,X_{k}\}$, such that the cardinality of $X^{r}$, denoted by $m$, is greater than that of $X^{e}$, denoted by $k$ ($|X^{r}|=m>|X^{e}|=k$).
@@
 $ITE^{\text{real}}=\alpha_{1}[u_{1}(x^{t}_{1})-u_{1}(x^{c}_{1})].$
-To compare the two causal effects without a salience effect, we establish a benchmark assumption that the salience ratio for any given dimension remains constant.
+To compare the two causal effects when alternatives have different attributes (and without a salience effect), we need a benchmark assumption that the salience ratio for any given dimension remains constant.
 Assumption 2.
@@
 For any non-zero $\alpha_{j}$, $\alpha_{k}$ and $\alpha^{\prime}_{j}$, $\alpha^{\prime}_{k}$ where $j,k\leq\min(m,n)$, the ratio $\frac{\alpha_{j}}{\alpha_{k}}=\frac{\alpha^{\prime}_{j}}{\alpha^{\prime}_{k}}$ holds.
-This assumption reflects the natural idea that if, in one circumstance, a DM considers attribute $j$ (e.g., gender) to be more salient than attribute $k$ (e.g., age), that is, $\alpha_{j}>\alpha_{k}$, then in another similar circumstance, the same DM should still perceive gender as more salient than age, $\alpha^{\prime}_{j}>\alpha^{\prime}_{k}$, even if the exact values of salience could differ ($\alpha_{j}\neq\alpha^{\prime}_{j}$ and $\alpha_{k}\neq\alpha^{\prime}_{k}$).
-The assumption implies that their relative magnitude should remain the same.
-This assumption is not related to the consistency of relative importance because the latter is defined on the causal effect rather than salience.
+This assumption reflects the natural idea that if, suppose a DM faces single one profile, DM considers attribute $j$ (e.g., gender) to be more salient than attribute $k$ (e.g., age), that is, $\alpha_{j}>\alpha_{k}$, then in another similar circumstance, again when DM faces only one profile, the same DM should still perceive gender as more salient than age, $\alpha^{\prime}_{j}>\alpha^{\prime}_{k}$, even if the exact values of salience could differ ($\alpha_{j}\neq\alpha^{\prime}_{j}$ and $\alpha_{k}\neq\alpha^{\prime}_{k}$).
+This holds true particularly because the attributes are uncorrelated and the DM evaluates only one profile at a time, eliminating any influence from other profiles.
+Thus, there is no reason for the order of salience to reverse.
+It is important to note that this assumption does not pertain to the consistency of relative importance, which is defined based on causal effects, not salience.
 Under this assumption, we can derive the relationship between $\alpha^{\prime}_{1}$ and $\alpha$, leading to the following result.
@@
 Specifically, following common practice, the experimental causal effect is measured by the AMCE.
-As noted in the SI 4, since AMCE is essentially a weighted sum of ITE, the amplification bias of ITE is also reflected in AMCE.
+As noted before, since AMCE is essentially a weighted sum of ITE, the amplification bias of ITE is also reflected in AMCE.
 2.1.1 Testing the Bias Amplification Hypothesis
@@
 3 Salience Effect
-Exploring further, we investigate the distortion in the linkage between experimental causal effects and real-world implications, moving beyond the assumption of fixed salience in the previous section.
+In this section, we explore the inconsistency beyond the assumption of fixed salience in the previous section.
 We adopt a psychological framework developed by Bordalo, Gennaioli and Shleifer (2012) to elucidate how variations in attribute salience can influence this relationship.
@@
 The critical condition within this proposition, “$\exists j$ and $k$, such that $r^{j}_{k}(x_{1})\neq r^{j}_{k}(\tilde{x}_{1})$,” elucidates a scenario where altering the treatment level affects the salience of at least one attribute within one profile.
-In the accompanying proof (detailed further in the SI 2), it is demonstrated that such a reversal can occur irrespective of the utility magnitudes associated with excluded attributes, or the inherent salience of the treatment attribute ($X_{1}$).
+In the proof (detailed further in the SI 2), it is demonstrated that such a reversal can occur irrespective of the utility magnitudes associated with excluded attributes, or the inherent salience of the treatment attribute ($X_{1}$).
 This finding has profound implications for experimental design and interpretation, particularly in contexts where the primary interest lies in understanding the direction of causal effects.
-It warns researchers of the possibility that, under the influence of salience effects, the observed direction may actually invert.
 To gain intuition regarding the phenomenon of Effect Sign Reversal, consider two hypothetical worlds, as depicted in the left panel of Table 2.
@@
 These two factors could lead to $ITE_{3}$ and $ITE_{2}$ having different signs.
-Based on Proposition 2, we formulate the following testable hypothesis:
+A weaker conclusion implied by this example and proposition 2 is that the effect observed in one scenario (the experimental world) may not necessarily reverse in a different scenario (the real world); instead, it may simply dissipate to no effect.
+Intuitively, if parameters exist so that they can support the case that pulls down a positive effect to be negative, it is even more likely that parameters could be found to support the case that reducing the effect to null.
+We formalize this observation in the following corollary:
+Corollary 2.
+Suppose assumptions 2 and 3 hold.
+$\exists j$ and $k$, such that $r^{j}_{k}(x_{1})\neq r^{j}_{k}(\tilde{x}_{1})$, then experimental treatment effect may become a null effect in the real world, and vice versa.
+Based on Proposition 2 and corollary 2, we formulate the following testable hypothesis:
 Hypothesis 2.
-The direction of attribute effects reverses with the different numbers of attributes in a conjoint experiment.
-We do not assert that an effect sign reversal will inevitably occur.
+The sign of the effect in a conjoint experiment may reverse or diminish to a null effect depending on the number of attributes involved.
+We do not assert that an effect sign reversal or diminish will inevitably occur.
 For instance, if one attribute—or its salience—predominantly overshadows others in the evaluation process, the reversal of this attribute’s effect is unlikely to be observed.
@@
 The direct consequence is that its sign will predominantly determine the sign of $V$, and thus the ITE.
-Another scenario that could preclude the observation of an effect sign reversal is the non-fulfillment of the rank change assumption.
+Another scenario that could preclude the observation of an effect sign reversal or diminish is the non-fulfillment of the rank change assumption.
 As highlighted by Proposition 2, the condition of salience rank change is critical for the observed result.
@@
 Hypothesis 3.
-Attribute effect direction reversal is less likely if the level of attributes changes only marginally.
-3.1.1 Testing the Effect Reversal Hypotheses
+Attribute effect direction reversal or diminish is less likely if the level of attributes changes only marginally.
+3.1.1 Test Hypotheses 2 and 3
 We utilize data from a conjoint experiment on hotel rooms conducted by Bansak et al.
@@
 Consequently, respondents were randomly assigned to one of 11 different experimental groups, which included profiles with 4, 5, 6, 7, 8, 9, 10, 12, 14, 18, or up to 22 attributes.
-Figure 4 presents a heatmap of the results (Figure A.3 in SI labels statistical significance).
+Figure 3 presents a heatmap of the results [^9] Figure A.3 in SI labels statistical significance.
+Due to low statistical power, we observe a limited number of significant estimates.
+As the sample size increases, we anticipate that the confidence intervals will narrow around the estimates, potentially leading to observations of significant effect reversals.
+However, observing a null effect is also consistent with our theoretical results and aligns with Hypothesis 2..
 The horizontal axis represents the number of attributes, and the vertical axis denotes the level of each attribute.
@@
 To facilitate a clearer visualization of the results, we assign each attribute a rank based on its absolute magnitude within each group.
-In the heatmap 5 (and (figure A.4 in SI), each column presents the relative importance rank of attributes in a specific experiment group, as listed on the x-axis.
+In the heatmap 4, each column presents the relative importance rank of attributes in a specific experiment group, as listed on the x-axis.
 The colors denote the importance rank.
+[^10] Figure A.4 in SI indicates statistical significance.
 Also, we construct importance by $R^{2}$ as Bansak et al.
@@
 Furthermore, our focus on Average Effects (AMCE) rather than Individual Effects (ITE) necessitates a particular consideration.
-As outlined in the SI 4, AMCEs are derived by marginalizing over attribute components and individuals, typically requiring participants to make several rounds of decisions among multiple profiles in a conjoint experiment, with each round featuring randomly generated profiles.
+As outlined in the SI 3, AMCEs are derived by marginalizing over attribute components and individuals, typically requiring participants to make several rounds of decisions among multiple profiles in a conjoint experiment, with each round featuring randomly generated profiles.
 If the randomly assigned attribute levels are uniquely distinct and lack counterparts in the real world, this can lead to distorted salience effects.
@@
 An immediate implication for research design is that researchers should endeavor to identify which attributes DMs truly value in real-life situations, thereby ensuring a match between the consideration sets in experiment and real-world contexts.
-As shown by Hainmueller, Hangartner and Yamamoto (2014), if the consideration sets are matched, survey experiments have ecological validity.
+As shown by Hainmueller, Hangartner and Yamamoto (2014), if the consideration sets are matched, survey experiments have consistency.
 This alignment requires a meticulous preliminary survey prior to the experimental design.
@@
 Such validation serves as a litmus test for the realism of the profiles and their alignment with real-world experiences.
-Pilot testing of the experiment with a select sample of participants from the target population would serve as a valuable step in this process.
-Feedback obtained from the pilot test could be used for fine-tuning the relevance and realism of the attributes and profiles.
-Based on this feedback, necessary adjustments can be made to the experimental design to ensure its close resemblance to real-world decision-making contexts.
-In the final stage of the experimental design process, the feedback from the pilot test and expert validation is integrated to refine the design.
-The culmination of this process is an experimental design that is not only aligned with real-world decision-making contexts but also equipped to provide meaningful insights into the dynamics of real-world decision-making processes.
 5 Concluding Remarks
@@
 Beyond the assurance of internal validity, external validity remains a longstanding concern.
-This article delves into a more fundamental aspect of generalizability: ecological validity.
-We theoretically and empirically demonstrate that experiments engaging in multi-dimensional decision-making processes, especially conjoint survey experiments, may lack ecological validity.
+This article delves into a more fundamental aspect of generalizability: the consistency of experimental and real-world effects.
+We theoretically and empirically demonstrate that experiments engaging in multi-dimensional decision-making processes, especially conjoint survey experiments, may lack such consistency.
 In particular, we find that the magnitude of experimental effects can be inflated, and the direction of effects and their relative importance may diverge from those observed in real-world scenarios.
-Through the lens of formal theory, this study uncovers how attention and salience effects could undermine ecological validity.
+A caveat to our findings is that we do not claim the inconsistency problem is inevitable in all such experiments.
+Many studies, such as those by Jenke et al.
+(2021), along with empirical examples discussed in our paper, suggest the stable consistency of certain attributes.
+Instead, we propose two plausible mechanisms and outline conditions that could lead to inconsistencies.
+From an empirical perspective, this encourages researchers to be more vigilant when designing experiments.
+Theoretically, we incorporate behavioral theories to deepen our understanding of experimental dynamics.
 The alignment between experimental conditions and real-world scenarios lays the groundwork for an in-depth exploration of external validity, involving variations in treatment, outcome measures, populations, and contexts.
-Developing a more comprehensive theory that integrates ecological validity with other forms of external validity represents a promising direction for future research.
+Developing a more comprehensive theory that integrates consistency with other forms of external validity represents a promising direction for future research.
 Supplementary Information
@@
 This study is pre-registered with the center for Open Science Framework (OSF).
-An anonymized version along with the questionnaire can be found at the OSF).
+An anonymized version along with the questionnaire can be found at the OSF.
 The survey experiment received full Institutional Review Board (IRB) approval in Study ID [details removed for anonymous peer review] from the authors’ institutions.
@@
 D.2 Pre-registered Design
-The online anonymized version along with the questionnaire can be found at the OSF).
+The online anonymized version along with the questionnaire can be found at the OSF.
 This study explores how the causal effects derived from conjoint experiments can be inconsistent with the real-world effect due to attention and salience biases.
@@
 (0.082) (0.087) (0.084)
-(1) (2) (3) (4) (5)
 RaceWhite 0.066 0.163 0.078
@@
 Children2 0.159
-(1) (2) (3) (4) (5)
 Children3 0.138
@@
 (0.108) (0.103) (0.116)
-(1) (2) (3) (4) (5)
 income32,000 --0.045 --0.338 0.021
@@
 child2 0.044
-(1) (2) (3) (4) (5)
 child3 0.158